### v2

# bCenter for Statistical Research and Methodology, U.S. Census Bureau, Suitland, MD

Abstract
Inference on overall ranking of a set of entities, such as athletes or players, schools and universities, hospitals, cities, restaurants, movies or books, companies, states, countries or subpopulations, based on appropriate characteristics or performances, is an important problem. Estimation of ranks based on point estimates of means does not account for the uncertainty in those estimates. Treating estimated ranks without any regard for uncertainty is problematic. We propose a novel solution using the Bayesian approach. It is easily implementable, competitive with a popular frequentist method, more effective and informative. Using suitable joint credible sets for entity means, we appropriately create credible distributions (CDs, a phrase we coin), which are probability distributions, for the rank vector of entities. As a byproduct, the supports of the CDs are credible sets for overall ranking. We evaluate our proposed procedure in terms of accuracy and stability using a number of applications and a simulation study. While the frequentist approach cannot utilize covariates, the proposed method handles them routinely to its benefit.
Keywords: Bayes; Credible sets; Fay-Herriot model; hierarchical Bayes; highest posterior density; small area estimation; unstructured Bayes.
1 Introduction
Reliable estimates of measures of important characteristics of agriculture, employment, health, education, for example, for various subpopulations are some of the important data produced by many national and international statistics offices around the world. For instance, the United Nations Sustainable Development Goals, the US Small Area Income and Poverty Estimation (SAIPE) and many programs in European Union need accurate poverty data at disaggregated levels for making public policies and implementing welfare programs. Disaggregated levels may be defined by partitioning a particular population into several subpopulations. For example, subpopulations may refer to the states of a country (population). While the term “subpopulations” has a wider meaning, we will also use “small areas” or “small domains” in its place. In a broader sense we use the word “entity” which may refer to a subpopulation or an individual.
Accurately estimating the means for one or more characteristics for the subpopulations is usually the primary goal of the government agencies. However, inference on overall ranking of entities, such as individuals, small areas, hospitals, herds of animals, etc. is also substantially important. Often it is also important to identify entities with very low or high means. Effective implementation of welfare or well-being programs overseen by authorities, identification of better yielding crop varieties in agriculture, or the top-breeding sires in animal breeding, etc. require reliable estimation of the rank vector. Limited resources mandate authorities to identify the underprivileged or impoverished subpopulations to provide assistance. Similarly, a business to open new stores/offices would like to know the localities with the most demands for its products and services.
These goals require accurately estimating the ranks of many or all subpopulations. According to Laird and Louis (1989), ranking draws attention to unusually good or poor performance of some subpopulations. They noted that ranking arises in a wide variety of applications, and appropriate solutions provide useful tools to investigators and policy makers for setting priorities for the next stage of their investigations.
The importance of joint ranking of $m$ subpopulations with means $\theta_{1},\cdots,\theta_{m}$ of a common characteristic has been emphasized in a pioneering article by Klein et al. (2020) (we will subsequently also use the abbreviation KWW to refer to this paper or the method by these authors). Overall true ranks of the subpopulations are unknown as those are determined by the true unknown $\theta_{i}$’s. Klein et al. (2020) noted that ranking of populations, governmental units, teams, players, health service providers, etc. based on experimental, observational or “survey data are usually released without direct statistical statements of uncertainty on estimated overall rankings”. They emphasized that prior to their work national statistical agencies did not adopt any statistical methodology to account for uncertainty of released estimates of overall rankings. They further cautioned that from published point estimates of the means only, practitioners frequently use those to naively rank the subpopulations. Such ranks are only point estimates, and ignore uncertainty in the estimates of $\theta_{i}$’s. Different sets of estimates of $\theta_{i}$’s from different data sets will potentially lead to different sets of ranks of the subpopulations. Clearly, even the best possible estimators of small area means are subject to error due to sampling variability of the estimators. It is thus both imperative and a sound policy to evaluate uncertainty associated with the reported ranks based on reasonable estimates of subpopulation means. Obviously, the accuracy of the estimated ranks can be appropriately studied through the joint distribution of the ranks, induced by the joint distribution of estimators of the means.
In an application to rank the fifty states of the U.S. and Washington, DC, using mean commuting times of workers who are sixteen or older and not working from home, Klein et al. (2020) used data from the American Community Survey (ACS). In their approach to the joint confidence set for the true rank vector, which they determine from an appropriate joint confidence set of the true means, the authors treated the unknown means non-random. They followed the classical approach where the unknown true parameters $\theta_{i}$’s are non-random. Some of the early frequentist works on ranking and selection of population means are by Bechhofer (1954) and Gupta (1956). While Klein et al. (2020) first created a joint confidence set for the mean vector $\theta=(\theta_{1},\cdots,\theta_{m})^{T}$ to create confidence set for ranks based on the frequentist approach, they briefly remarked that an analogous Bayesian version of their frequentist solution is possible through construction of a joint credible set for $\theta$. Our Bayesian solution is not based on what Klein et al. (2020) alluded to in that remark. Our solution is novel and to implement it we need a Bayesian paradigm. We have proposed here two easily implementable solutions to this problem.
In our Bayesian approach, we first obtain an empirical credible set for $\theta$ by selecting a subset of $\theta$ values from an MCMC sample that are in a suitable region. Subsequently, we use these selected values to create probability distributions of the components of the rank vector. To convey these probability distributions originate from a credible region we coined the phrase “credible distributions” (CDs) to refer to these distributions. Our proposed solutions admit better interpretations and appear to be more accurate based on several desirable measures (see Sections 6 to 10). They are at least as good as, if not better than, the frequentist solution. A reviewer brought it to our attention to a very recent paper by Mogstad et al. (2024) on frequentist approach to overall ranking. These authors extended the KWW work by considering joint confidence intervals of all pairwise differences of the form $\theta_{i}-\theta_{j}$. As in the approach by KWW, the confidence intervals of the differences of the means are used to create joint confidence estimation of overall ranking. Since our objective is to develop an intuitive and easily implementable Bayesian solution to the overall ranking, we stick here only to the seminal work by Klein et al. (2020) for comparison with our Bayesian solution.
The Bayesian approach treats $\theta$ as random and assigns them a suitable joint distribution. Following the Bayesian approach to ranking several binomial populations by Bland and Bratcher (1968), Govindarajulu and Harvey (1974) presented a unified Bayesian treatment of the ranking and selection problem. Goel and Rubin (1977) proposed a Bayesian approach on selecting a subset containing the best subpopulation. Laird and Louis (1989) used an empirical Bayes (EB) approach to the ranking problem to environmental monitoring of regions suspected of elevated disease rates for residents or health service providers suspected of elevated services failure rates for patients, requiring attention to mitigate risks. Indeed, Morris and Christiansen (1996) highlighted the utility of Bayesian approach to the ranking of hospitals. Unlike the frequentist approach of Klein et al. (2020), the EB and the (fully) Bayesian approaches facilitate hierarchical modeling of the random latent means $\theta$ through regression on relevant available covariates. An early application of hierarchical modeling to ranking of subpopulations is due to Efron and Morris (1975). In their application to rank thirty-six cities of El Salvador based on prevalence of toxoplasmosis rates, they indicated the pitfall of ranking using only observed estimates of the rates but ignoring the variability of these estimates.
In Section 2 we provide a brief review of the relevant literature on ranking. We introduce two reasonable Bayesian models by specifying suitable joint prior distributions for the unknown subpopulation means. Corresponding to a prior distribution a Bayesian model is formed by augmenting the sampling distribution of the data, which is the cornerstone of the classical approach to a frequentist solution. In our Bayesian formulation, we consider two noninformative improper prior distributions. The prior distribution in equation (4) that provides no structure for the unknown $\theta_{i}$’s is justified since joint credible sets for the $\theta_{i}$’s from the posterior distribution exactly agree with the joint confidence sets of the $\theta_{i}$’s developed by Klein et al. (2020). The second prior distribution in Section 2.2 is based on a conjugate prior for the means that leads to a widely used hierarchical model, introduced earlier by Fay and Herriot (1979), Berger (1985), Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989) and Morris and Christiansen (1996). While most of these authors used an EB approach by estimating the parameters of the prior distribution in the hierarchical model, Berger and Deely (1988) and Morris and Christiansen (1996) used the hierarchical Bayes (HB) approach (or its approximation) by assigning reasonable noninformative, minimally influential, improper priors to the model parameters.
We follow these latter groups and pursue a fully Bayesian (that is, HB) approach with similar, minimally influential improper priors on the model parameters. We add that whereas Berger and Deely (1988) and Morris and Christiansen (1996) used the HB approach for (point) estimation of individual ranks, we consider this approach for joint set estimation of the rank vector. Our solution, elaborated in Section 4, is based on empirical representation of a $100(1-\alpha)\%$ credible set for $\theta$ using MCMC samples of $\theta$ from the posterior distribution. The MCMC samples that are in the credible region are taken to represent the empirical credible region (ECR). Each of these samples in the ECR is ranked, and in totality they induce relative frequency distributions, defined earlier as credible distributions, of the ranks of the entities.
Here is a brief outline of the paper. In Section 2 we introduce some useful notation. In Subsections 2.1 and 2.2 we present two Bayesian models that are the building blocks of our solution. In Section 3 we present a brief review of the Klein et al. (2020) frequentist solution. In Section 4 we describe empirical credible sets. In Subsection 4.1 we discuss empirical Cartesian and elliptical (related to highest posterior density) credible sets for $\theta$. In Section 5 we describe computation of credible distributions (CDs). In Section 6 we compare various set estimators (confidence and credible) of $\theta$ based on suitable size measures. In Section 7 we introduce two new measures to compare the frequentist and Bayesian solutions to the overall ranking problem. In Section 8 we apply and contrast our solution to the Klein et al. (2020) solution for ranking of the US states based on commuting times. In Section 9 we develop credible distributions to rank the U.S. states based on median incomes. In Section 10 we develop credible distributions for the baseball data studied by Efron and Morris (1975). In Subsection 10.1 we present simulation results under the baseball example setup to compare various solutions for overall ranking. Section 11 provides a summary and a few concluding remarks. In Subsection S.2 we review some results on volume and derive average representative length of an elliptical set. In Subsection S.4 we present a full discussion on the baseball example.
2 A selected review of the ranking literature and introduction of Bayesian models
Our motivation for this research is to develop a Bayesian solution to the joint ranking problem proposed in a pioneering paper by Klein et al. (2020). Suppose $\theta_{i}$ is the mean for certain characteristic of the $i$th subpopulation, $i=1,\cdots,m$. In their application, Klein et al. (2020) are interested in ranking the fifty states of US and Washington, DC, based on $\theta_{i}$, the mean commuting times of the workers not working from home, $i=1,\cdots,m=51$. Similar to theirs, our goal is also to make joint inference about the true ranks $\check{r}_{1},\cdots,\check{r}_{m}$ (see equation (1)), of the $m$ subpopulations based on estimators $Y_{1},\cdots,Y_{m}$ of $\theta_{1},\cdots,\theta_{m}$, and appropriate models. Alternatively, these subpopulations may be called entities. In ranking $\theta_{i}$’s we treat them continuous variables. We rank the $\theta_{i}$’s from the smallest to the largest, where the rank of $\theta_{i}$ is
$\check{r}_{i}\equiv\check{r}_{i}(\theta)=\sum_{j=1}^{m}I(\theta_{j}\leq\theta_{i})=1+\sum_{j:j\neq i}I(\theta_{j}\leq\theta_{i}),~{}\mbox{for }i=1,\cdots,m.$ (1)
Based on $Y_{1},\cdots,Y_{m}$, estimators of $\check{r}_{1},\cdots,\check{r}_{m}$ are $R_{1},\cdots,R_{m}$, where
$R_{i}=\sum_{j=1}^{m}I(Y_{j}\leq Y_{i})=1+\sum_{j:j\neq i}I(Y_{j}\leq Y_{i}),~{}\mbox{for }i=1,\cdots,m.$ (2)
Obviously, the estimators $R_{i}$’s are subject to random errors, which stem from the point estimators $Y_{i}$’s. It is important to account for uncertainty in the estimators $R_{i}$’s. Klein et al. (2020) proposed a clever solution to account for uncertainty in the joint estimation of the true rank vector. While they adopted a frequentist approach, we will present purely Bayesian solutions as alternatives based on two Bayesian models for the $\theta_{i}$’s. Actually, to estimate individual ranks $\check{r}_{i}$’s of the $\theta_{i}$’s, Aitkin and Longford (1986) considered setups where in one case they assumed $\theta_{i}$’s are non-random, and in the second case they modeled $\theta_{i}$’s by some probability distribution, possibly depending on parameters ($\psi$, say). The latter case constitutes a hierarchical modeling for the $Y_{i}$’s and $\theta_{i}$’s, and estimates of $\theta_{i}$’s or $\check{r}_{i}$’s can be determined by an EB approach if $\psi$ is estimated, or a fully Bayesian (FB) approach if $\psi$ is also assigned a prior distribution, proper or improper. Aitkin and Longford (1986) followed the EB approach for hierarchical model. Berger and Deely (1988) proposed a fully Bayesian approach to test a null hypothesis ($H_{0}$) for equality of subpopulation means, and conditional on $H_{0}$ is false, they estimated probability of $\theta_{i}$ being the largest of the $m$. They estimate the ranks of the $\theta_{i}$’s based on the ranks of these probabilities.
In the remainder of the article, for compactness of notation, we will often use $Y,y,\theta$ and $\check{r}$ to denote the vectors $(Y_{1},\cdots,Y_{m})^{T},(y_{1},\cdots,y_{m})^{T},(\theta_{1},\cdots,\theta_{m})^{T}$ and $(\check{r}_{1},\cdots,\check{r}_{m})^{T}$, respectively. Here, $y$ is the observed value of the random vector $Y$. Klein et al. (2020) assumed for the estimators $Y_{1},\cdots,Y_{m}$, that
$Y_{i}|\theta\;\stackrel{{\scriptstyle ind}}{{\sim}}\;N(\theta_{i},D_{i}),~{}i=1,\cdots,m,$ (3)
where the variances $D_{i}$ associated with $Y_{i}$ are assumed to be known. In the small area estimation literature $D_{i}$’s are often referred to as the sampling variances.
To account for uncertainty of the rank estimators $R_{1},\cdots,R_{m}$, the authors first constructed, for some specified $\alpha\in(0,1)$, a $(1-\alpha)$-level joint confidence set for $\theta$ based on confidence intervals $I_{i}=[Y_{i}-z_{1-\frac{\gamma}{2}}\sqrt{D_{i}},Y_{i}-z_{\frac{\gamma}{2}}\sqrt{D_{i}}]\equiv[L_{i},U_{i}],i=1,\cdots,m,$ for $\theta_{i}$’s, where $z_{q}$ is the $q$th quantile of the standard normal distribution. They determined $\gamma$ from $\alpha$ in two ways: by applying (i) Bonferroni’s inequality ($\gamma=\alpha/m$), and (ii) independence of the $Y_{i}$’s, so that the joint confidence level is $(1-\alpha)(=(1-\gamma)^{m})$. Using the individual intervals, they formed the Cartesian product $I_{1}\times I_{2}\cdots\times I_{m}$ to obtain the joint confidence set of the true mean vector $\theta$. They used this confidence set to develop a $(1-\alpha)$-level joint confidence set for the true rank vector $\check{r}$. We briefly review their method in Section 3 below. Bonferroni’s method does not need independence of the pivotal quantities to create confidence set for $\theta$ and it is broadly applicable. However, the intervals are usually slightly wider than those produced by the independent method. This may result in additional overlapping intervals and a slightly more expansive set for overall ranking (cf. the rank sets for these two methods for HI and TX, in Table 2 of KWW).
While the frequentist joint confidence set developed by Klein et al. (2020) accounts for the uncertainty in the estimated rank vector $(R_{1},\cdots,R_{m})$, some limitations to their method are that (i) it does not provide any opportunity to utilize any relevant covariate that may be related to the $\theta_{i}$’s, and (ii) the generated set is overly “stretched out”. As mentioned earlier, following Berger and Deely (1988) we will assign to the $\theta_{i}$’s a similar probability model and present a fully Bayesian solution for a joint set estimation of the overall ranking. We emphasize that although Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989) and Morris and Christiansen (1996) considered EB/HB estimation of individual ranks based on hierarchical modeling, none considered the (credible) set estimation of the rank vector $(\check{r}_{1},\cdots,\check{r}_{m})^{T}$. Our proposed solution, defined and explained in Section 5, is based on credible distributions.
2.1 An unstructured Bayesian (UB) model
Here, we augment the sampling model in equation (3) by the prior pdf for $\theta_{1},\cdots,\theta_{m}$
$\pi(\theta_{1},\cdots,\theta_{m})=1~{}\mbox{for}~{}-\infty<\theta_{1}<\infty,\cdots,-\infty<\theta_{m}<\infty,$ (4)
which is an improper and noninformative prior. The joint posterior pdf of $\theta_{1},\cdots,\theta_{m}$ is
$\pi(\theta_{1},\cdots,\theta_{m}|y_{1},\cdots,y_{m})=\prod_{i=1}^{m}N(\theta_{i}|y_{i},D_{i}),$ (5)
where $N(\theta_{i}|y_{i},D_{i})$ denotes the pdf of a normal distribution with mean $y_{i}$ and variance $D_{i}$. The main reason for the noninformative uniform prior for $\theta_{1},\cdots,\theta_{m}$ over the $m$-dimensional Euclidean space in equation (4) follows from the fact that a reasonable (in particular, HPD, highest posterior density) $1-\gamma$ credible interval for $\theta_{i}$ is identical to confidence interval $I_{i}$ for $\theta_{i}$ that was the cornerstone in the Klein et al. (2020) solution. If we denote the Cartesian product, $I_{1}\times\cdots\times I_{m}$ of $I_{i}$’s by $\Theta_{C,u}$, then it will yield a joint credible set for $\theta$ that is identical to the joint confidence set constructed by Klein et al. (2020).
We will use this credible set to create one set of credible distributions solutions for the rank vector $\check{r}$. While the unstructured Bayes model helps us to reproduce the frequentist confidence set of KWW, this model fails to adequately address the impact of one or more large sampling variances $D_{i}$ on the credible set of the rank vector $\check{r}$. Indeed, if $y_{i}>y_{j}$, then, irrespective of their sampling variances, $P[\theta_{i}>\theta_{j}|y]>1/2$. In other words, if $y_{1}$ is the largest of the $y$-values, then even if its sampling variance $D_{1}$ is the largest and excessively large, with a posterior probability more than 0.5, $\theta_{1}$ will exceed any other $\theta_{i}$. In particular, we noted in our applications below, that in such cases the posterior probability of $\check{r}_{1}=m$ turned out to be more than the probability that $\check{r}_{i}=m$ for any $i\neq 1$. Berger and Deely (1988) and Morris and Christiansen (1996) justifiably criticized this result as unsatisfactory.
2.2 A hierarchical Bayesian (HB ) model
To address the perceived weakness of the UB model, we augment model (3) by specifying a suitable model for $\theta_{1},\cdots,\theta_{m}$. This corresponds to a popular hierarchical model:
  • (I)
    $Y_{i}|\theta_{1},\cdots,\theta_{m}\stackrel{{\scriptstyle ind}}{{\sim}}N(\theta_{i},D_{i}),i=1,\cdots,m,$
  • (II)
    $\theta_{i}|\beta,A\stackrel{{\scriptstyle ind}}{{\sim}}N(x_{i}^{T}\beta,A),i=1,\cdots,m$,
where $\beta$, a $p\times 1$ regression coefficient, and $A$ the model variance, are the model parameters. The $p\times 1$ known vector $x_{i}$ includes covariates associated with the $i$th subpopulation. This model was used by many, including Efron and Morris (1975), Berger (1985), Berger and Deely (1988) and Morris and Christiansen (1996). This model was first used by Fay and Herriot (1979) for EB in small area estimation (SAE). In the SAE literature this is the celebrated Fay-Herriot model that is extensively used for EBLUP and EB approaches to SAE, e.g., p. 77 of Rao and Molina (2015).
The step (II) above, links $\theta_{i}$ to $x_{i}$, provides a structure for the means $\theta_{i}$’s through the covariate $x_{i}$. It is known as the linking model in SAE, see again Rao and Molina (2015). If $x_{i}=1$, it corresponds to an intercept only model, which results in an exchangeable structure for the $\theta_{i}$’s. If the model error variance $A$ is small, then the covariates $x_{i}$’s introduce a lot of structure to the $\theta_{i}$’s. On the other hand, if $A$ is large, there will be little structure for the $\theta_{i}$’s. Indeed, if $A\rightarrow\infty$, the model (II) does not control any structure for the $\theta_{i}$’s through $x_{i}$’s, it would be identical to the UB model for $\theta_{i}$’s in Section 2.1.
For known model parameters $\beta,A$, the Bayes estimators of $\theta_{i}$’s are obtained from the posterior pdf of $\theta_{i}$’s, which are independent normal distributions. The posterior pdf’s depend on $\beta$ and $A$. Efron and Morris (1975) suggested estimate the parameters from the marginal pdf of $Y_{1},\cdots,Y_{m}$ (obtained from the joint pdf of $Y_{1},\cdots,Y_{m},\theta_{1},\cdots,\theta_{m}$ from (I) and (II) above), and use the estimated (based on estimates of model parameters) posterior pdf to make inference about $\theta_{i}$’s. The resulting solution is known as the EB solution.
On the other hand, Berger (1985), Berger and Deely (1988), and many others, pursued a fully Bayesian approach, the HB approach, by specifying a prior pdf, typically, noninformative and improper, for $\beta,A.$ In our applications, we use the improper prior pdf
$\pi(\beta,A)={(\bar{D}+A)^{-\frac{1}{2}}},$ (6)
where $\bar{D}=\sum_{i=1}^{m}D_{i}/m$. While we use $\bar{D}+A$ to define our prior pdf, Berger and Deely (1988) used the geometric mean of $(D_{i}+A)$’s to define their pdf.
Aitkin and Longford (1986), Laird and Louis (1989) and Morris and Christiansen (1996) used EB or HB estimates of $\theta_{i}$’s in place of $Y_{i}$’s for estimating the individual ranks $\check{r}_{i}$’s and probability that $\theta_{i}$ is the largest of the means. The EB solutions to these problems have explicit expressions. Berger and Deely (1988) showed that HB solution depended on evaluating several one-dimensional integrals with respect to $A$. The hierarchical model in (I) and (II), complemented by a prior pdf on $\beta,A$, as in equation (6), is known as the Fay-Herriot HB model in the SAE literature.
3 A brief review of joint confidence set of overall ranking by Klein et al. (2020)
To review Klein et al. (2020) solution for joint confidence set of overall ranking, we start with the confidence interval $[L_{i},U_{i}]$ for $\theta_{i}$, based on $Y_{i}$, for $i=1,\cdots,m$. These intervals are such that the (frequentist) probability $P\{\cap_{i=1}^{m}[L_{i}\leq\theta_{i}\leq U_{i}]\}$ is no less than $(1-\alpha)$. For each $i\in\{1,2,\cdots,m\}={\cal M}$, they define various subsets of ${\cal M}$, given below,
$C_{i}$ $=$ $\{1,2,\cdots,m\}\setminus\{i\},~{}\Lambda_{Li}=\{j\in C_{i}:U_{j}\leq L_{i}\},~{}\Lambda_{Ri}=\{j\in C_{i}:U_{i}\leq L_{j}\},$
$\Lambda_{Oi}$ $=$ $\{j\in C_{i}:U_{j}>L_{i}~{}\mbox{and}~{}U_{i}>L_{j}\}=C_{i}\setminus\{\Lambda_{Li}\cup\Lambda_{Ri}\}.$ (7)
Using these subsets, in Section 3 Klein et al. (2020) show that the set of rank vectors
$\{(r_{1},\cdots,r_{m}):r_{i}\in\{|\Lambda_{Li}|+1,|\Lambda_{Li}|+2,|\Lambda_{Li}|+3,\cdots,|\Lambda_{Li}|+|\Lambda_{Oi}|+1\}~{}\mbox{for}~{}i=1,\cdots,m\},$ (8)
is a joint confidence set for the overall rank vector $\check{r}=(\check{r}_{1},\cdots,\check{r}_{m})$ having a joint coverage probability at least $(1-\alpha)$. For a set $G\subset{\cal M}$, the number of elements in $G$ is $|G|$. The joint confidence set for the vector $\check{r}$ is specified by the sets $\Lambda_{Li},\Lambda_{Oi},i=1,\cdots,m$, which depend on the joint confidence sets of $\theta_{i}$’s, that are determined by the estimators $Y_{i}$’s and their measures of uncertainty.
In the frequentist approach above, the ranking is indirect through the relative positioning of individual confidence intervals. This is unlike the credible distributions (CDs) under the Bayesian approach we shortly propose where samples of $\theta$ are generated from the posterior distribution and the ranking is direct based on $\theta$ samples. The frequentist algorithm above relies very much on comparison of confidence intervals of components of $\theta$. It requires that a joint confidence set must be an orthotope (hyperrectangular) so that each component of $\theta$ can vary freely without being constrained by the other components. To its disadvantage, the algorithm does not allow use of more efficient joint confidence regions, for example, likelihood-based regions, since ranges of $\theta_{i}$’s in these regions are usually interdependent.
4 Description of empirical credible sets
For our Bayesian models we create credible distributions by a sampling-based approach. In this approach we generate a large sample of size $S$ of $\theta_{1},\cdots,\theta_{m}$ values from their joint posterior pdf, $\pi_{M}(\theta|y)$, derived under the model $M$, where $M$ is $UB$ for unstructured, and $HB$ for hierarchical model. Generically, we denote these posterior sample values of $\theta=(\theta_{1},\cdots,\theta_{m})^{T}$ by the set $\{\theta^{(1)},\cdots,\theta^{(S)}\}$, where $\theta^{(s)}$ is an $m\times 1$ vector.
For the UB model in Subsection 2.1, any posterior sample $\theta^{(s)}$ is a vector of $m$ independent normal variables with mean $y_{i}$ and variance $D_{i}$, $i=1,\cdots,m$. For the HB model in Subsection 2.2, we draw these samples via MCMC, which in our applications is straightforward. See Subsection S.1 in the Supplement for details.
A point estimator $\hat{\theta}$ by itself does not usually provide any uncertainty (for example, measure of variability) associated with it. Any reasonable confidence set (or credible set, in Bayesian approach) for $\theta$ is an acceptable way to account for variability of $\hat{\theta}$. Usually, an appropriately constructed confidence set based on the distribution of $\hat{\theta}$ will be “tight” if the variability of $\hat{\theta}$ is small and vice versa.
Based on any point estimator $\hat{\theta}$ we can get a point estimator $\widehat{\check{r}}$ of the rank vector $\check{r}$. Just as the estimator $\hat{\theta}$ by itself does not provide an estimator of variability, the estimator $\widehat{\check{r}}$ also does not provide any measure of variability. If $\hat{\theta}$ is accurate, so will usually be the estimator $\widehat{\check{r}}$. If $\widehat{\check{r}}$ is not accurate, it may be much different from the true rank vector $\check{r}$. Thus, it makes sense to use a better estimator of $\theta$, for example, model-based EB estimator $\hat{\theta}_{EB}$ or HB estimator $\hat{\theta}_{HB}$, than the direct estimator $Y$. But any point estimator of rank $\check{r}$ based on a model-based estimator of $\theta$ still does not account for uncertainty.
Estimators of individual ranks based on model-based estimators of $\theta$ have been considered by Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989), Goldstein and Spiegelhalter (1996) or Morris and Christiansen (1996). Like ours, the approach to this problem by Berger and Deely (1988) and Goldstein and Spiegelhalter (1996) is fully Bayesian. The last group of authors considered a Bayesian approach to comparison of several health or educational institutions. They developed for their applications 95% credible intervals for individual ranks. However, none of these estimation methods considers joint overall ranking. Klein et al. (2020) noted that assessment of the uncertainty in the estimated overall ranking would involve all the populations simultaneously and their relative standing to each other.
For our Bayesian solutions we follow KWW only to construct a suitable joint credible set for $\theta$. Subsequently, our approach to creating set estimate for overall ranking vector is much different and, evidence shows, our solution is more informative. To explain our ideas in simple terms, we consider $\theta_{1}$, which is a scalar. Based on a large number of sampled values $\theta_{1}^{(1)},\cdots,\theta_{1}^{(S)}$ from the posterior distribution of $\theta_{1}$, their $(\delta/2)$th and $(1-\delta/2)$th quantiles reasonably provide an empirical $1-\delta$ credible interval for $\theta_{1}$. Intuitively, all the $\theta_{1}^{(s)}$ values which are inside this empirical credible interval may be considered “plausible values” of $\theta_{1}$. By extending this idea analogously to the whole vector $\theta$, first we draw ${\cal F}_{M}=\{\theta^{(1)},\cdots,\theta^{(S)}\}$, a sample from the joint posterior distribution $\pi_{M}(\theta|y)$ of $\theta$. Next we construct empirically a suitable joint $(1-\alpha)$ credible set, ${\cal F}_{M,\alpha,y}$, for $\theta$ based on the samples ${\cal F}_{M}$. We present construction of two intuitively appropriate credible sets below. It is reasonable to consider that the elements of ${\cal F}_{M,\alpha,y}$ are plausible values of $\theta$. We will use these plausible values to create credible distributions (explained below) for the ranks $\check{r}_{1},\cdots,\check{r}_{m}$.
4.1 Constructing ${\cal F}_{M,\alpha,y}$: an empirical joint credible set for $\theta$
For some $\alpha\in(0,1)$, we determine a suitable subset of ${\cal F}_{M}$, ${\cal F}_{M,\alpha,y}$ so that the set has approximately $S(1-\alpha)$ elements. This set is an empirical $(1-\alpha)$ joint credible set of $\theta$ under the model $M$.
4.1.1 A Cartesian empirical credible set for $\theta$
This approach closely follows only the confidence set construction by Klein et al. (2020) for $\theta$; our credible solution for $\check{r}$ is different. From the vectors in ${\cal F}_{M}$, for $i=1,\cdots,m$, we take the $i$th component of each vector to create a sample ${\cal F}_{i}$ for $\theta_{i}$ from its marginal posterior distribution. For some suitable $\kappa\in(0,1)$, we determine the $(\kappa/2)$th and $(1-\kappa/2)$th quantiles, $a_{i}$ and $b_{i}$, respectively, of the set ${\cal F}_{i}$. Let ${\cal S}_{i}=\{s:a_{i}\leq\theta_{i}^{(s)}\leq b_{i},s=1,2,\cdots S\}$. Define the set ${\cal S}_{J}=\cap_{i=1}^{m}{\cal S}_{i}$. Let $K_{J}=|{\cal S}_{J}|$. For any $\alpha\in(0,1)$, we will determine $\kappa$ in such a way that $K_{J}$ is the integer closest to $S(1-\alpha)$. For this choice of $\kappa$, the set $\{\theta^{(s)},s\in{\cal S}_{J}\}$ is an empirical $(1-\alpha)$ joint credible set of $\theta$. For a given $\alpha$, an initial choice of $\kappa$ is $1-(1-\alpha)^{1/m};$ however, we need to iteratively modify $\kappa$ to satisfy the approximate relation $K_{J}\approx S(1-\alpha)$ as closely as possible. Finally, corresponding to the selected set $\{\theta^{(s)},s\in{\cal S}_{J}\}$, we obtain an approximate $(1-\alpha)$ joint credible distribution for the rank vector $\check{r}$.
In Remark 8 of their paper, Klein et al. (2020) suggested constructing a Cartesian credible set for $\theta$ and using this set to create a credible set for the overall ranking by applying the main result of their paper. We note that the joint credible set for $\theta$ constructed this way by these authors for the Bayesian model in Section 2.1 will be identical to their frequentist confidence set, but this time the result will be interpreted conditional on the data instead of conditional on the $\theta$ as is done in the frequentist case. However, we show by simulations and a real application to median income estimation (in Section 9) that our Bayesian solution to ranking produces a smaller error in estimation measured by expected posterior absolute deviation from the “true ranking,” or some “gold standard” for it.
4.1.2 An elliptical empirical credible set for $\theta$
An optimal method to create credible set for $\theta$ is the highest posterior density (HPD) method. A point $\theta_{*}$ in the parameter space of $\theta$ is eligible for inclusion in an HPD credible set $\Theta_{H}$ if the posterior density $\pi_{M}(\theta_{*}|y_{1},\cdots,y_{m})$ is no less than the density at any point outside the set $\Theta_{H}$. We continue adding such $\theta_{*}$ values to the HPD set until the credible level of $\Theta_{H}$ is at least $(1-\alpha)$. The method is intuitively sound since it includes the more likely $\theta$ values in terms of the posterior distribution. Obviously, an HPD credible set is of the form $\{\theta:\pi_{M}(\theta|y_{1},\cdots,y_{m})\geq t\}$, where the positive threshold $t$ depends on $y,\alpha$, and is chosen by satisfying the credible level $(1-\alpha)$. In terms of Lebesgue volume, among all $(1-\alpha)$ credible sets, the HPD credible set has the lowest volume.
For the UB model in Section 2.1, for a constant $c$ an HPD credible set is of the form
$\left\{\theta:\sum_{i=1}^{m}(y_{i}-\theta_{i})^{2}/D_{i}\leq c\right\}.$ (9)
The HPD credible set above, elliptic shape, shows that if a point $\theta$ is inside the set, the components of that point cannot vary freely since a larger value of $(y_{i}-\theta_{i})^{2}$ will make a larger value of $(y_{j}-\theta_{j})^{2}$ for $j\neq i$ less likely. This is in direct contrast to the Cartesian credible region in Subsection 4.1.1. It should be mentioned that for the frequentist approach it is possible to get an elliptical confidence set of the form in (9) by inverting an appropriate likelihood-based acceptance region. However, due to interdependence of values of the components of $\theta$ that are allowed in the region, it would not be possible to line up these components to determine their relative rankings. This is a major hurdle with the use of the optimal likelihood-based confidence region to determine confidence region for overall ranking $\check{r}$. However, shape of a credible region poses no challenge to the ranking problem in our Bayesian setup. Since we can deal with the values of $\theta$ inside a credible set directly, determining the overall ranking of the components is not a problem. It is a major advantage to the Bayesian approach.
The credible set in equation (9) depends on the Mahalanobis distance (MD) function
$M(\theta,y,D)=(y-\theta)^{T}D^{-1}(y-\theta),$ (10)
where, for the UB model, $y$ is the posterior mean and $D=\mbox{diag}(D_{1},\cdots,D_{m})$ is the posterior variance-covariance matrix of $\theta$. Based on a posterior sample ${\cal F}_{UB}$ of $\theta$, we can compute the set of MD values $\{M(\theta,y,D):\theta\in{\cal F}_{UB}\}$ and take the $(1-\alpha)$ quantile of these values for the cut-off $c$, and construct an empirical HPD credible set given by
${\cal F}_{UB,\alpha,y}=\{\theta\in{\cal F}_{UB}:M(\theta,y,D)\leq c\}.$
For the more general HB model in Section 2.2, for large $m$ the posterior pdf $\pi_{HB}(\theta|y)$ is approximately multivariate normal (it follows heuristically by noting that the conditional posterior distribution of $\theta$ given $A$ is a multivariate normal distribution, and applying the Laplace approximation and consistency of the posterior mode of $A$). Thus, an HPD credible set for $\theta$ will be approximately elliptic, defined by the Mahalanobis distance function based on the posterior mean and variance-covariance of $\theta$. Specifically, we can construct an empirical HPD credible set using the sample of $\theta$ values, ${\cal F}_{HB}$, from the posterior pdf. To this end we compute estimated posterior mean and variance-covariance matrix, given by
$\hat{\theta}_{HB}=\frac{1}{S}\sum_{s=1}^{S}\theta^{(s)},~{}~{}V_{HB}=\frac{1}{S}\sum_{s=1}^{S}\{\theta^{(s)}-\hat{\theta}_{HB}\}\{\theta^{(s)}-\hat{\theta}_{HB}\}^{T},$
respectively. Next, we compute the Mahalanobis distances $d^{(s)}_{HB}=M(\theta^{(s)},\hat{\theta}_{HB},V_{HB}),$ $s=1,\cdots,S$. With $c=(1-\alpha)$-quantile of these distances, we create an empirical credible set ${\cal F}_{HB,\alpha,y}$ for $\theta$, similar to ${\cal F}_{UB,\alpha,y}$, based on $d^{(s)}_{HB}$ values. This set is a sample-based empirical representation of the HPD credible set for $\theta$ with credible coefficient $(1-\alpha)$. Note that $c\approx\chi^{2}(1-\alpha,m)$, the $(1-\alpha)$-quantile of chi-square distribution with $m$ degrees of freedom.
5 Joint distribution of overall ranking on a $(1-\alpha)$ credible set: credible distributions
We have underscored the importance of accounting for uncertainty for a set of released ranks based on some point estimates of the subpopulation means. We now present a simple and straightforward solution using a Bayesian approach. Bayesian inference for any function, possibly vector-valued, of the parameter can be conducted by computing a multi-dimensional relative frequency distribution of the function based on a reasonably large sample of $\theta$ from its posterior distribution.
From a sample $\theta^{(1)},\cdots,\theta^{(S)}$ from $\pi_{M}(\theta|y)$ we select $\theta^{(s_{1})},\cdots,\theta^{(s_{K})}$, to be denoted by ${\cal F}_{M,\alpha,y}$. For each sample in ${\cal F}_{M,\alpha,y}$, we then create an $m\times m$ two-way table, columns as the entities to be ranked and rows as the ranks of the entities. For the sample $\theta^{(s_{1})}$, we start with an $m\times m$ null matrix. If the $i$th component of $\theta^{(s_{1})}$ does not tie with any other component and has rank $j_{i}$, we put a $1$ in the $j_{i}$th row of the $i$th column. If two components $k$ and $l$ tie for ranks $j$ and $j+1$, we replace elements $(k,j),(k,j+1),(l,j),(l,j+1)$ by $1/2$. We proceed similarly with ties among more than two components. In this way for all $K$ values of $\theta$ in ${\cal F}_{M,\alpha,y}$, we complete $K$ tables, and we take a weighted average of these tables, where the weights sum to $K$ so that the resulting table has its each row and column sum to $K$. Two possible choices of the weights are: (a) equal weights, or (b) proportional to $\exp[-M(\theta^{(s)},\cdot,\cdot)/2]$, where $M(\theta^{(s)},\cdot,\cdot)$ is the suitable Mahalanobis distance. Throughout this paper, the “unweighted” method actually corresponds to equal weights (case (a) above).
By converting the entries of this table to fractions so that all rows and all columns sum to 1, we obtain a double stochastic square matrix, which plays a pivotal role in this paper. Its $k$th row assigns probabilities to various entities to hold the rank $k$. The $i$th column of this matrix is providing an empirical posterior distribution of rank ${\check{r}}_{i}$, the rank of the $i$th entity. We name this probability distribution as the credible distribution (CD). For later use we write this probability distribution of ${\check{r}}_{i}$ by $\pi_{{\check{r}}_{i},M,T,y}$. We added the subscript “T” to indicate the type of credible sets: T=C for Cartesian, and T=E for elliptical. With respect to this distribution, for a function $g({\check{r}}_{i})$, we use $E^{\pi_{{\check{r}}_{i},M,T,y}}[g({\check{r}}_{i})]$ to denote its posterior expectation.
6 Comparison of set estimators of $\theta$ by size
Two types of set estimators for the mean vector $\theta$ are playing the critical part in accounting for uncertainty in the inference for overall ranking. Klein et al. (2020) used a confidence region for the mean vector which is the Cartesian product of appropriate confidence intervals of the component means. In our Bayesian effort to closely resemble the frequentist set for $\theta$, we have also developed a credible set, which is an orthotope or a hyperrectangle, a Cartesian product of suitable credible intervals for the individual means. As an alternative to this Cartesian set, where each $\theta_{i}$ varies freely, we have also proposed HPD sets for $\theta$. It is intuitively clear that in terms of reasonable “size” measures HPD sets would be more “compact” than sets of other shapes. In this section, we use two obvious measures of set estimators of $\theta$ with comparable confidence or credibility level, and compare various sets.
Two size measures that appear to be appropriate for a set estimator is its Lebesgue volume and average length of the “sides” determining the set. A Cartesian set estimator $C$ is given by $\prod_{i=1}^{m}[L_{i},U_{i}]$. Its volume, $V_{C}$, and average length, $L_{C}$, respectively, are
$V_{C}=\prod_{i=1}^{m}(U_{i}-L_{i}),~{}~{}L_{C}=\frac{1}{m}\sum_{i=1}^{m}(U_{i}-L_{i}).$ (11)
The other credible regions we have considered are elliptical in shape, generically defined by
$E=\{\theta:(\theta-\theta^{*})^{T}\Phi(\theta-\theta^{*})\leq c\},$ (12)
where $\theta^{*}$ is an appropriate center, $\Phi$ is an appropriate $m\times m$, p.d. matrix, and $c>0$ is a suitable cut-off, all known. The Lebesgue volume $V_{E}(\Phi,m,c)$ of the set $E$ is well known, and its expression is shown in equation (S.2) in the Appendix. A suitable length measure of an elliptical region is not available. We have attempted a reasonable length measure in the Appendix. The length measure, denoted $L_{E}$, is
$L_{E}=\frac{1}{m}\sum_{i=1}^{m}L_{M,i},$ (13)
where the “length” $L_{M,i}$ is provided by equation (S.5) in the Appendix.
For HPD or approximate HPD credible sets, the matrix $\Phi$ in equation (12) corresponds to $V_{HB}^{-1}$ in HB models and $\Phi=D^{-1}$ in the UB model.
7 Comparison of frequentist and Bayes solutions to the rank inference problem
Frequentist and Bayes solutions to the set inference problem for ranks are very different. The frequentist KWW solution is a Cartesian product of subsets of $\{1,2,\cdots,m\}$. The Bayesian solution is through probability distributions, the CDs, on $\{1,2,\cdots,m\}$. Given their different nature, these solutions can be compared only indirectly.
7.1 Comparison via entropy
A KWW confidence interval of $\check{r}_{i}$, rank of entity $i$, is an interval $\{a_{i}+1,\cdots,a_{i}+n_{i}\}$ for a positive integer $n_{i}$ and a nonnegative integer $a_{i}$. A CD of $\check{r}_{i}$ is a probability distribution on $1,\cdots,m$. These solutions are not directly comparable. In the absence of any other information for a KWW set, we need to assume the possible values of $\check{r}_{i}$ are uniformly distributed on the KWW set, and calculate entropy or other deviation measure that would compare the “concentrations” provided by these solutions. The entropy of a probability distribution measures its uncertainty or concentration. The entropy of a discrete distribution $f$ is
$E=-\sum_{x:f(x)>0}f(x)\log(f(x)),$
see, for example, p.91, Berger (1985). Lower entropy means a distribution is more concentrated or informative. The lowest entropy is 00, attained only by a degenerate distribution. We apply this concept to our applications.
7.2 Comparison via cardinality
To evaluate a deviation measure mentioned above, we can proceed as follows. Again, for a discrete random variable $W$ taking values $a_{i}+1,\cdots,a_{i}+n_{i}$, each with probability $1/n_{i}$, it easily follows that $4E\Big{(}|W-\mu|\Big{)}=n_{i}-n_{i}^{-1}C$, where $C$ is zero if $n_{i}$ is even, and $C=1$ if $n_{i}$ is odd, where $\mu=E(W)$. An importance of this result is that if the KWW rank set for $\check{r}_{i}$ contains $n_{i}$ points $\{a_{i}+1,\cdots,a_{i}+n_{i}\}$, then $n_{i}\approx 4E|W-\mu|$.
8 Ranking of the US states based on commuting times
In a pioneering article that motivated our study, Klein et al. (2020) applied their frequentist approach to rank fifty states of the U.S. and DC by mean commuting times of workers sixteen or older and not working from home. They used survey data collected from the ACS. In Figure 1, we recreated Figure 1 of Klein et al. (2020) that depicted the frequentist solution of the confidence region for ranking from equation (8). Possible ranks for state $i$ in this set are shown by a pink line segment stretching from $|\Lambda_{Li}|+1$ to $|\Lambda_{Li}|+|\Lambda_{Oi}|+1$. From the figure, for example, the possible ranks for the state ID are $4-9$. On the other hand, the states which can hold the rank 9 are the states WY, AK, MT, IA, KS and ID (the pink line segments for these states intersect the horizontal line for Ranks = 9). The rank based on direct estimate $y_{i}$ is shown by a black dot at its rank $r$. If two states tie for ranks $r,r+1$, their tied ranks are shown by a black line segment joining these states at rank $r+1$. See, for example, SD and ND direct estimates tie for ranks 1 and 2.
Figure 1: An alternative visualization of the 90% joint confidence region for travel time ranking from Figure 1 of KWW with overlapping CDs from UB Cartesian sets (weighted)
We overlaid on Figure 1 the CDs based on a Cartesian credible set from the UB model. Both in terms of the model and the method, this table is most directly comparable with Figure 1 of Klein et al. (2020). Probabilities from CDs are shown by yellow circles, bigger circles for larger probabilities.
To clearly interpret the probabilities depicted in the figure, let us focus specifically on its $4$th row and $4$th column. The 4th column represents the state of NE has ranks $3-6$ with probabilities $0.21$, $0.20$, $0.55$ and $0.04$, respectively. We can contrast this with the red band on the 4th column in the figure which shows the rank set for NE suggested by the Klein et al. (2020) solution. The band shows that in addition to the four ranks $3-6$, NE can also rank 1, 2, 7 and 8.
On the other hand, the fourth row of the figure corresponding to rank 4 shows that the states that can plausibly hold this rank are AK, MT, NE and WY (with respective probabilities $0.04$, $0.23$, $0.55$ and $0.18$). Again, in contrast, the 4th row in Figure 1 of Klein et al. (2020) shows that states SD, ND, WY, NE, MT, AK, IA, KS and ID are all contenders of rank 4. Compared to our Bayesian CD solution, the frequentist solution includes five additional states. However, our CD for ID shows that it is virtually impossible for ID to be a contender of rank 4.
Again, if we consider ID, we find from Figure 1 that ID can hold rank 9 with probability virtually 1, and the rank 9 can only be held by ID. However, we observed it earlier that for the frequentist solution possible ranks for ID are $4-9$, and the states that may be ranked 9th are WY, MT, AK, IA, KS and ID. Similarly, while SD and ND, tied by direct estimates, both have the same frequentist rank set $1-6$, whereas in terms of Bayesian CDs, only probable ranks for these states are $1$ and $2$, and SD (ND) has $60\%\ (40\%)$ probability for rank $1$. Clearly, all these demonstrate that our Bayesian results are “less dispersed”, “more illuminating” and “more conclusive”.
On a related note, we observe that based on the facts that $y_{50}=y_{51}$ and they are the smallest among all $y$’s and $D_{51}<D_{50}$, it can be shown that for the HB model, $P(\check{r}_{51}>\check{r}_{50}|y)=P(\theta_{51}>\theta_{50}|y)<1/2$. Thus, we may take it as a justification that under the CDs $\{\check{r}_{51}=1\}$ is more likely than $\{\check{r}_{50}=1\}$; based on the HB CDs, we got these probabilities $\approx 0.6$ vs. $\approx 0.4$.
Table 1: Sizes CD-s-HB-C of (weighted) Cartesian HB CDs, KWW-c cardinality of KWW sets, entropies CD-e-HB-C of (weighted) Cartesian HB CDs, and KWW-e of KWW sets for ranks of nine states that have lower values of direct estimates of commuting times
States ID KS IA AK MT NE WY ND SD Average
CD-s-HB-C 1.00 2.23 1.93 3.84 3.46 2.31 4.49 1.95 1.95 2.57
KWW-c 6 7 7 9 9 8 9 6 6 7.44
CD-e-HB-C 0.01 0.74 0.88 1.39 1.33 1.12 1.42 0.69 0.68 0.92
KWW-e 1.79 1.95 1.95 2.20 2.20 2.08 2.20 1.79 1.79 1.99
We compare in Table 1 the KWW frequentist solution with the HB Cartesian CDs based on their “sizes” and “informativeness” using a subset of 9 states that have lower values of direct estimates of commuting times. First two rows measure the size and show that the CDs have smaller “size”. Last two lines measure the informativeness which show that the CDs are more “informative”.
For these 9 states, we also compute the credible distributions (i.e., the double stochastic matrix introduced in Section 5), shown in Table S.1 in Supplement, based on the elliptical region from the HB model. From the $4$th row of this matrix we find that the states AK, MT, NE and WY are the contenders of rank 4, and they may hold this rank with probabilities $0.11,0.21,0.50$ and $0.18$, respectively. The fourth column of that matrix conveys that NE can hold the ranks 3, 4, 5 and 6 with probabilities $0.32,0.50,0.15$ and $0.04$, respectively. Agreement between this set of probable values and those for the other cases such as the UB Cartesian show that credible regions of $\theta$ having different shapes and derived from different models provide reasonably similar solutions for overall ranking. From Table S.1 it is clear that with practical certainty that ID has rank 9, AK can have ranks $3-8$, and SD can have ranks 1 or 2. The corresponding KWW confidence sets, respectively, are $\{4-9\},\{1-9\}$ and $\{1-6\}$. This solution would assert that the ranks 4 for ID, 5 for AK and 6 for SD are plausible. However, this assertion would be strongly contradicted by the double scatter plot in Figure S.1. The plot makes it abundantly clear that commuting time for ID is higher than for AK, which in turn is higher than that for SD.
We have no covariate for this example and the HB estimate of the model error variance $A$, $13.2$ is very large; when put in perspective it is about $56$ times the largest sampling variance $D_{i}$ and $3572$ times the smallest one. With so high ratios of the model variance to sampling variances, the HB model is practically identical to the UB model. Similarity of the CDs from the UB and the HB models in this application are confirmed by the simulation study we presented in Table 4. In the last block of the table we set $A=1$ which is nearly $188$ times the largest sampling variance and $342$ times the smallest one. We find from the two lower-right-bottom blocks of this table that the average posterior mean absolute deviations reported for the UB and the HB columns show no practical differences among the four (HB and UB models for each of the Cartesian and elliptical credible sets) Bayesian solutions.
We also compare all five (one frequentist and four Bayesian) confidence and credible regions of $\theta$ in terms of their size measures. From Table 2 we find that all Bayesian credible sets except the UB-Cartesian one have smaller volumes than the KWW confidence set. While volumes of Cartesian credible sets are similar to the volume of the KWW set, the elliptical shape HPD credible sets from both models are enormously (by an order of $10^{-7}$) superior (this is due to cutting off low probability “corners” of the hyperrectangular sets to create an HPD set). Even in terms of average length, elliptical sets are better than the hyperrectangular sets. Although the KWW confidence intervals and UB Cartesian credible intervals are the same for $\theta$, when measuring the “sizes” of the KWW sets and the CDs for ranks, the average sizes are significantly smaller for the CDs (see the last line of Table 2).
Table 2: Commuting time: sizes of frequentist and Bayes set estimators of $\theta$ and $\check{r}$.
Frequentist Unstructured Bayes (weighted) Hierarchical Bayes (weighted)
Parameter Measure KWW Cartesian Elliptical Cartesian Elliptical
$\theta$ Volume 4.52 4.52 $1.66\times 10^{-7}$ 4.38 $1.56\times 10^{-7}$
$\theta$ Average length 1.16 1.16 0.83 1.16 0.83
$\check{r}$ Ave cardinality of rank sets 10.02 2.12 2.12 2.39 2.39
9 Ranking of the US states based on median incomes
Inference for median incomes of the US states for the income year 1989 has served as a benchmark example to evaluate effectiveness of various small area estimation methods. See for example, Ghosh et al. (1996) and Chung and Datta (2022). The Current Population Survey (CPS) provides annual estimates of median incomes for the states for various family sizes. These estimates serve as the direct estimates for the popular Fay-Herriot model. The application enables borrowing strength from other covariates by integrating administrative data and past census data. For this application two useful covariates are available. The first covariate, $x_{i1}$, is the $i$th state median income for 1979 from the 1980 Census. The second covariate, $x_{i2}$, is an adjustment to $x_{i1}$ for the $i$th state, $x_{i2}=(\text{PCI}_{i,1989}/\text{PCI}_{i,1979})x_{i1}$, $i=1,\ldots,m$, where $\text{PCI}_{i,1979}$ and $\text{PCI}_{i,1989}$ are the 1979 and 1989 per capita incomes of the $i$th state provided by the Bureau of Economic Analysis of the U.S. Department of Commerce. The 1989 census incomes for the states, which we denote by $\omega_{i}$, were medians based on a large number of households from each state. Due to large sample sizes many researchers (see, for example, Ghosh et al. (1996) and Chung and Datta (2022)) considered $\omega_{i}$’s as “gold standards” for true incomes, $\theta_{i}$, which we are trying to rank. We rank these surrogate values and treat their ranks, $\xi_{i}$, to be the surrogates for the true ranks $\check{r}_{i}$.
Figure 2: A visualization of the 90% joint confidence region for median income ranking by KWW with an overlapping credible distribution from the Cartesian HB method (weighted)
In Figure 2, we recreated Figure 1 of KWW that depicted the frequentist confidence region for ranking from equation (8). On this figure we overlaid the Cartesian credible distribution based on the weighted HB method. The KWW rank sets for all the states are depicted by appropriate yellow line segments that span a subset of 1 to 51. Probabilities from CDs are shown by red circles, bigger circles for larger probabilities. We use solid black circles on this figure to depict the “true” states ranks based on median incomes from the 1990 census. To the state CT the KWW solution assigns ranks 15 to 51, which is very wide. In contrast, the HB Cartesian CD assigns probabilities $0.03,0.13,0.33,$ and $0.51$, respectively to the ranks 48 to 51. For NJ the KWW ranks span from 22 to 51. The HB solution assigns probabilities $0.04,0.53,$ and $0.43$, respectively, to the ranks $49-51$. Clearly, the Bayesian CDs are very compact and informative compared to the frequentist solution.
Table 3: Median incomes: Comparison of set estimators of $\theta$ and $\check{r}$, derived by frequentist and Bayesian approaches (all weighted), by various measures
Parameter Measures $\epsilon_{i,\small{HB,C}}$ $\epsilon_{i,HB,E}$ $\epsilon_{i,UB,C}$ $\epsilon_{i,UB,E}$ $\epsilon_{i,KWW}$
$\check{r}$ Total abs. dev. 179.52 179.52 388.11 388.11 798.66
$\check{r}$ Ave entropy 2.27 2.27 2.62 2.62 3.86
$\check{r}$ Ave cardinality 10.27 10.27 15.57 15.57 47.94
$\theta$ Volume/$g$ $4.619\times 10^{7}$ $2.975$ $8.321\times 10^{14}$ $4.043\times 10^{7}$ $1.080\times 10^{15}$
[…]
In Table 3 we compare the frequentist and Bayesian solutions based on various measures. Line 1 compares the KWW solutions and the CDs using respective deviations from the surrogate values of ranks from the 1990 census based on equation (14). Line 2 compares these solutions in terms of their average entropies. Line 3 compares the solutions based on “sizes” of the rank solutions. Lines 4 and 5 list the scaled volumes and scaled ave lengths of the confidence/credible sets of $\theta$. Volumes of Bayesian sets (of $\theta$) are vastly smaller than those of the frequentist confidence sets. The HB elliptical set has the smallest volume. Average lengths of Bayesian sets are smaller than the average length of confidence set, again the HB elliptical set has the smallest average length.
Coming to line 1 of Table 3, we use the aforementioned surrogate values $\xi_{i}$ of the ranks $\check{r}_{i}$ to assess accuracy of various overall rankings. In particular, we compute for entity $i=1,\cdots,m,$
$\epsilon_{i,HB,C}$ $=$ $E^{\pi_{{\check{r}}_{i},HB,C,y}}[|\check{r}_{i}-\xi_{i}||y],~{}\epsilon_{i,HB,E}=E^{\pi_{{\check{r}}_{i},HB,E,y}}[|\check{r}_{i}-\xi_{i}||y],$
$\epsilon_{i,UB,C}$ $=$ $E^{\pi_{{\check{r}}_{i},UB,C,y}}[|\check{r}_{i}-\xi_{i}||y],~{}\epsilon_{i,UB,E}=E^{\pi_{{\check{r}}_{i},UB,E,y}}[|\check{r}_{i}-\xi_{i}||y],$
$\epsilon_{i,KWW}$ $=$ $\frac{1}{|\Lambda_{Oi}|+1}\sum_{j=|\Lambda_{Li}|+1}^{|\Lambda_{Li}|+|\Lambda_{Oi}|+1}|j-\xi_{i}|,$ (14)
where $\epsilon_{i,HB,C}$ is the posterior expectation of the absolute deviation of $\check{r}_{i}$ from $\xi_{i}$ with respect to the credible distribution $\pi_{{\check{r}}_{i},HB,C,y}$ from the HB model, and similarly, the other expectations. Finally, $\epsilon_{i,KWW}$ is the average of absolute deviations of $\xi_{i}$ from the possible values for $\check{r}_{i}$ given by the KWW confidence set $\{|\Lambda_{Li}|+1,\cdots,|\Lambda_{Li}|+|\Lambda_{Oi}|+1\}$ for $\check{r}_{i}$.
In Table S.2 we presented for each state average absolute deviations (cf. equation (14)) of the KWW ranks from the surrogate ranks based on 1989 income data from the 1990 census. We also presented the posterior expected absolute deviations from (14) for the credible distributions (for the elliptic version only). Deviations for the KWW method exceed those for the HB method for all states, and $46$ KWW deviations except OK, GA, CO, NV, and DC exceed those for the UB method. We find that the KWW method has the highest sum of average absolute deviations (799), about 344% larger than for the HB method, and about 116% larger than for the UB method. Clearly, the best method is HB. The table shows that in terms of total of component expected absolute deviations of the rank solutions from the surrogate ranks (from equation (14), the CDs are vastly superior to the KWW solution. The HB CDs are better than the UB CDs but the shape of the credible sets for $\theta$ appears not to play any role. Similar conclusions emerge from the other two measures in Table 3 for evaluating solutions for $\check{r}$.
For the median income application we further explore comparison of credible distributions generated by various Bayesian models (UB and HB, with and without covariates) and methods (weighted and unweighted) among themselves and with the frequentist method of KWW in terms of various deviations of equation (14). We plotted several of these posterior expected absolute deviations against the KWW average absolute deviations in Figure 3.
The left panel in the first row of Figure 3 is the plot using posterior expected absolute deviations from credible distributions from weighted method based on the HB model with covariates on the $y$-axis and the KWW deviations on the $x$-axis. We use green circles for the elliptical method and black stars for the rectangular method. This panel shows that all the points cluster below the $y=x$ line and closer to the $x$-axis, indicating that the credible distributions are vastly superior to the KWW method. It also shows that for the weighted method there are practically no differences between the deviations from credible distributions based on the rectangular or the elliptical credible sets for the mean vector $\theta$.
The right panel in the first row of that figure differs from the left panel only by use of no covariates in the HB model. We see that several points are above the $y=x$ line, indicating that the KWW solution for those states are better than the HB solution in the absence of covariates. Also, although many of these points that are below the $y=x$ line, a large number of points have $y$-values bigger than 5, in contrast with the left panel where only about $15\%$ of the points are that big, indicating that the credible distributions without covariates are not as good as those created by using good covariates. These two panels convincingly demonstrate the importance of effective covariates for the ranking problem.
The left panel on the second row mimics the left panel on the first row except that now the credible distributions are unweighted. The HB model employs covariates. As we have noticed in the first row, in this case also we see that the Bayesian solutions (both elliptical and rectangular) are vastly better than the KWW solution. However, while most of the points on the panel in the second row have $y$-values bigger than 5, most are below that in the panel on the first row. These two panels show the superiority of the weighted credible distributions. In the right panel on this row we plotted the deviations based on weighted, elliptical credible distributions for the UB model and the HB model without covariates against the KWW deviations. While overall the credible distributions show superiority over the KWW method, for about $10\%$ of the states the KWW method performed better. None of these methods utilized any covariates. This plot also shows that in the absence of a covariate the HB method loses its advantage over the UB method, and the two credible distributions performed rather similarly.
In the last row of Figure 3, we compare various CDs (weighted vs. unweighted, elliptical vs. Cartesian) from the UB model against the confidence set provided by the KWW method in terms of posterior expected absolute deviations and average absolute deviations resulting from the KWW method (computed from (14)). While again the UB credible distributions show overall superiority over the KWW method, for about $10\%$ of the states the KWW method performed better. None of these methods utilized any covariates.
Figure 3: Various plots of expected or average absolute deviations for median incomes
Refer to caption Refer to caption
With Covariates Covariates not used
Refer to caption Refer to caption
With Covariates No covariates, elliptical
Refer to caption Refer to caption
Covariates cannot be used Covariates cannot be used
10 Ranking of baseball players: An example by Efron and Morris (1975)
Efron and Morris (1975) considered batting averages of 18 major league baseball batters from their first 45 at bats in the 1970 baseball season to predict their performances in the remainder of that season. Let $\theta_{i}$ be the unknown true batting average of player $i$ during that season. A reasonable representative value for $\theta_{i}$ was the player’s average ($\omega_{i}$) in the remainder of the season. The sample proportion of hits $Y_{i}$ from first 45 at bats is approximately normal with mean $\theta_{i}$ and variance $D_{i}$. In our analysis, we use the estimated variance, given by $D_{i}=Y_{i}(1-Y_{i})/45.$ Treating $\omega_{i}$ as the “gold standard” value of $\theta_{i}$, the total empirical squared error (TESE), $\sum_{i=1}^{18}(y_{i}-\omega_{i})^{2}$, for the direct estimates $y_{i}$’s, is $0.07537$, whereas $0.02368$ for the HB, and $0.02356$ for the empirical best linear unbiased predictor (EBLUP). The TESE for the direct estimates is about $3.2$ times of that of HB or the EBLUP estimates.
Efron and Morris (1975) used the arc-sin transformed parameters $\psi_{i}=arcsin(2\theta_{i}-1)$, for the transformed data $\hat{\psi}_{i}=arcsin(2Y_{i}-1)$ in their EB analysis. This transformation resulted in equal approximate sampling variances, $1/45$, for all $\hat{\psi}_{i}$’s, $i=1,\cdots,18$. They applied Stein’s shrinkage estimation method based on an exchangeable model to calculate ${\hat{\psi}}_{J-S,i}$, the James-Stein estimate of $\psi_{i}$. The TESE of prediction for the direct estimates $\hat{\psi}_{i}$ from $arcsin(2\omega_{i}-1)$ is 0.3902, which is about 3.5 times of 0.1113, the TESE for the James-Stein estimates ${\hat{\psi}}_{J-S,i}$. This application showed a substantial efficiency of Stein’s method. The efficiency for HB estimators of $\psi_{i}$ based on the same model is about $3.4$, which is comparable. Note that these numbers compare well with $3.2$, we obtained above for EBLUP/HB predictors of $\theta_{i}$’s.
Figure 4: A visualization of the KWW 90% joint confidence region for overall ranking of baseball players with overlapping CDs for UB and HB models (weighted, elliptical method)
By equation (9) of Klein et al. (2020) a $90\%$ joint confidence region for the rank vector $\check{r}$ is given by
$\{(\check{r}_{1},\cdots,\check{r}_{18}):\check{r}_{k}\in\{1,2,\cdots,18\}\mbox{ for }k=1,\cdots,18\}.$
This confidence set for the rank vector is depicted in Figure 4. Possible ranks for each player in this set are shown by a yellow line segment stretching from 1 to 18. According to this confidence set, any player can rank from 1 to 18, and any particular rank can be held by any of these 18 players; clearly, results are totally uninformative. Due to multiple testing, the critical value $2.76$ from equation (13) of Klein et al. (2020) is too large to declare any comparison between two direct estimates to be significant. A joint $90\%$ credible distributions for the rank vector is a double-stochastic (DS) matrix. This DS matrix is pictorially presented by overlaying on the figure of the confidence set solid red circles with area proportional to the probability it is representing. The top matrix is based on the elliptic set for $\theta$ under the UB model and the bottom matrix is a similar version for the intercept only HB model (both weighted).
From the top panel (UB) of Figure 4, it is clear that many probabilities in this matrix are practically zeros implying that those combinations of players and ranks are improbable. Undoubtedly, the resulting credible distributions are lot more informative than the frequentist confidence set. The bottom panel of this figure, corresponding to the HB intercept only, again shows that many red circles are still tiny, meaning that those combinations are less likely. However, in comparison with the top panel, these circles, or the probabilities, are bigger. This means that the CDs in the intercept case are less concentrated. This happens due to shift of the posterior distribution of each true mean $\theta_{i}$ to a common grand value (estimate of the intercept). Due to space limitation, a full discussion of this example is moved to Supplemental Section S.4.
In this example using the surrogate ranks $\xi_{i}$’s for the players, we computed the posterior expected absolute deviation of $\check{r}_{i}$ from $\xi_{i}$ using equation (14) for both the UB and the HB models. These values for all players are displayed in Table S.3. Thirteen of the eighteen deviations for the UB model-based solution are smaller than those for the HB model. Total of these deviations for the UB model are also smaller than its HB counterpart. The HB model performs worse. In contrast, the shrinkage estimates of $\theta$ (HB estimates based on the intercept model) produced a far less TESE, $0.02368$, compared with $0.07537$, the TESE of the UB model-based estimates of $\theta_{i}$’s. In the absence of any covariate, the HB model-based estimation of $\theta$ and ranks $\check{r}_{i}$’s go in the opposite directions. Similar results were obtained by Goldstein and Spiegelhalter (1996) (see their Tables 6b and 7b). But if good covariates are available that effectively explain the variability of $\theta$ (implied by small model error variance $A$ in the Fay-Herriot model), the credible distributions from HB models perform better than the CDs from an UB model. Simulation results that we report in Subsection 10.1 illustrate this point. A careful reading of the first two rows in each of the first three blocks for $A=0.001,0.005$ and $0.10$ in Table 4 clearly show this point. These results also shed light on why the CDs from the HB model in the median income example do significantly better than those from the UB model (cf. Figure 3 in Section 9).
10.1 A simulation comparing set estimators of overall ranking
To run a simulation for a setup similar to the baseball example we generate 18 $x_{i}$’s for a covariate independently from a uniform (0,1) distribution. These covariates are kept fixed throughout the simulation. Using them, we generate $\theta_{i}$’s independently from $N(\beta_{0}+\beta_{1}x_{i},A)$. Using the generated $\theta_{i}$, we generate $Y_{i}$ from $N(\theta_{i},D_{i})$. We use the $D_{i}$ values from the baseball application. We use five possible values of $A=0.001,0.005,0.01,0.1$ and $1$ (for the baseball data estimate of $A$ is about $0.002$). We use $\beta_{1}=0$ (no covariate) and $\beta_{1}=0.4$. We use $\beta_{0}=0.2$, although $\beta_{0}$ does not influence the ranks and the results.
We report our simulation results in Table 4. First row of each block of the table, with a W in parentheses, reports average posterior expected deviations for weighted credible distribution, and the second row with UW in parentheses, shows results for its unweighted version. In the left half of these tables we presented results with a covariate and in the right half with no covariate. We also provided the average of the mean absolute deviations of the overall ranks from the true ranks for KWW method, and average of the posterior expected absolute deviations for Bayesian methods: HB and UB. In the last two rows of each block, we presented average volumes and average of lengths of the sides of various 90% confidence or credible sets of the true mean vector $\theta$.
Table 4: Summary of simulations for the baseball setup
With covariate Without covariate
H. Bayes U. Bayes H. Bayes U. Bayes
$A$ KWW Cart Ellip Cart Ellip KWW Cart Ellip Cart Ellip
Ave Post Exp Abs Dev (W) 5.518 1.178 1.178 2.313 2.313 5.978 5.366 5.366 4.642 4.642
Ave Post Exp Abs Dev (UW) 5.518 1.543 1.521 2.560 2.545 5.978 5.194 5.199 4.804 4.795
0.001 (Ave Volumes)1/m 0.356 0.262 0.212 0.357 0.286 0.356 0.259 0.209 0.357 0.287
Ave length of the sides 0.357 0.247 0.199 0.358 0.287 0.357 0.241 0.196 0.358 0.287
Ave Post Exp Abs Dev (W) 5.268 1.835 1.835 2.148 2.148 5.919 4.065 4.066 3.468 3.468
Ave Post Exp Abs Dev (UW) 5.268 2.033 2.020 2.378 2.365 5.919 4.094 4.094 3.714 3.701
0.005 (Ave Volumes)1/m 0.356 0.297 0.238 0.357 0.287 0.356 0.295 0.237 0.356 0.286
Ave length of the sides 0.357 0.289 0.232 0.358 0.287 0.357 0.281 0.227 0.358 0.287
Ave Post Exp Abs Dev (W) 4.879 1.874 1.874 1.980 1.980 5.668 2.967 2.967 2.766 2.766
Ave Post Exp Abs Dev (UW) 4.879 2.063 2.053 2.188 2.176 5.668 3.211 3.203 3.013 2.999
0.010 (Ave Volumes)1/m 0.356 0.318 0.255 0.357 0.286 0.356 0.317 0.254 0.357 0.287
Ave length of the sides 0.357 0.314 0.252 0.358 0.287 0.357 0.312 0.251 0.358 0.288
Ave Post Exp Abs Dev (W) 2.919 1.070 1.070 1.061 1.061 3.212 1.177 1.177 1.168 1.168
Ave Post Exp Abs Dev (UW) 2.919 1.205 1.198 1.194 1.186 3.212 1.328 1.319 1.313 1.305
0.100 (Ave Volumes)1/m 0.356 0.351 0.282 0.356 0.287 0.356 0.350 0.281 0.356 0.286
Ave length of the sides 0.357 0.352 0.283 0.358 0.287 0.357 0.351 0.282 0.358 0.287
Ave Post Exp Abs Dev (W) 1.209 0.417 0.417 0.418 0.418 1.181 0.370 0.370 0.369 0.369
Ave Post Exp Abs Dev (UW) 1.209 0.467 0.465 0.467 0.464 1.181 0.426 0.423 0.426 0.423
1.000 (Ave Volumes)1/m 0.356 0.356 0.286 0.357 0.287 0.356 0.356 0.286 0.357 0.287
Ave length of the sides 0.357 0.357 0.287 0.358 0.288 0.357 0.357 0.287 0.358 0.288
When compared in terms of size measures, the $m$th root of the volume (to avoid dealing with very large or very small numbers) and the average length of the confidence or credible sets of $\theta$, the HB elliptical credible set performed the best. In terms of the average posterior expected deviations of ranks from their true values, the performance of the HB solution is mixed and the Klein et al. (2020) method has the weakest performance. Smaller values of $A$ (0.001, 0.005, 0.01), relative to the $D_{i}$’s, introduce a lot of structure to the $\theta_{i}$’s. In these cases in the absence of any covariate, $\theta_{i}$’s tend to be a lot similar, where the UB method performs better based on deviation criterion. However, when the $\theta_{i}$’s vary according to a covariate, the HB method performs substantially better for smaller values of $A$. For larger values of $A$, $0.1$ and $1$, even with the covariate, $\theta_{i}$’s do not display much of any structure (relative to the $D_{i}$’s, $A=0.1$ is between 19 and 35 times the $D_{i}$’s, and $A=1$ is between 190 and 350 times larger). The case of $A=1$ relative to the $D_{i}$’s is similar to our example where we develop ranking of the US states based on average commuting times. For the last two cases of $A$, the HB posterior distribution of $\theta$ will have negligible shrinkage and it is approximately equal to UB posterior distribution, resulting in similar CDs for overall ranking; the UB method is marginally better than the HB method in terms of the deviation measure. Finally, in terms of this measure, the overall ranking for the elliptical method does better than the Cartesian method when the double stochastic matrix is created by unweighted method. However, when the CDs are created by weighted method, the CDs from elliptical and the Cartesian sets perform very similarly.
11 Summary and Conclusions
Inference for population means (or proportions) of subpopulations or subjects is an important statistical problem. Estimation of (i) small area characteristics in survey sampling, (ii) breeding values of various herds in animal science and livestock research, (iii) true performance of various players in sports, and (iv) treatment effectiveness for certain conditions treated by several health facilities in healthcare, for example, are some of the important problems in multi-population studies. The major focus in statistical inference for multiple populations has been estimation (point and interval) of means or proportions or similar measures from several subpopulations or individuals. In small area estimation, the companion problem of simultaneously estimating the ranks of the small areas is also of considerable importance but it has not gotten enough attention. Only Shen and Louis (1998) simultaneously considered estimation of small area means and ranks. Using the HB approach these authors considered point estimation of ranks but did not account for error in estimation of the rank vector for all small areas.
To account for estimation error for the rank vector of many subpopulations, in a pioneering article Klein et al. (2020) considered set estimation of overall ranking. Klein et al. (2020) created joint confidence set of the rank vector by employing a rectangular confidence set of the mean vector $\theta$. Their solution in equation (8) attributes $|\Lambda_{Li}|+1,\cdots,|\Lambda_{Li}|+|\Lambda_{Oi}|+1|$ as the possible values for the true rank ${\check{r}}_{i},i=1,\cdots,m.$ This solution, frequentist by construction, may be interpreted as a confidence distribution for ${\check{r}}_{i}$, uniformly distributed over the integers between the lower and the upper bounds presented above. As an alternative to this method, which cannot use auxiliary information or model the unknown true mean $\theta$, we introduced a Bayesian model for the mean vector. The model allows us to utilize useful auxiliary information as well as to develop various credible sets for $\theta$ including the HPD sets that are exact or nearly exact. We created various credible distributions for the true rank ${\check{r}}_{i}$, which turn out to be more informative than the confidence set by KWW. In our Bayesian proposal we can also allow more probable $\theta$’s (measured by the posterior pdf) in the credible set to have them bigger weights to the credible distributions.
Application of proposed CDs is computationally straightforward. We applied our solution to three applications, two to rank small areas (discussed extensively) and one to rank professional baseball players (reported briefly accompanied by extensive discussion in the Supplement). All three applications and a simulation study mimicking two of the applications demonstrated superior performance of the proposed CDs over the existing frequentist method.
For the baseball example there are no reliable covariates but there are surrogate values of the true means from the players’ performance from the remainder of the season. Evaluation of frequentist confidence set and the credible distributions based on the average or expected absolute deviations of estimated ranks from the ranks of the surrogate values show better performance for the Bayesian solutions. In the absence of covariates, the UB credible distributions perform better than the HB credible distributions. For the commuting time example we neither have any covariates nor any representative values of the unknown true means. Even then in this example the credible distributions are “more conclusive”, “less dispersed” and “more illuminating” than the frequentist solution by Klein et al. (2020). For the median income estimation problem we have useful covariates as well as representative values (from the 1990 census) of the true means. Credible distributions from the HB model with the covariates performed the best when measured in terms of average absolute deviations. Both credible distributions, from the UB model and the HB model, have better performance than the frequentist method. The HB credible distributions based on covariates are uniformly better than the other two solutions. However, the HB credible distributions without any covariates do not enjoy a uniform superiority over the other two solutions. These outcomes are confirmed by our simulation study for the baseball setup. Availability of good covariates shows the superiority of HB credible distributions. Lack of covariates degrades the performance of HB credible distributions, which are sometimes inferior to the UB credible distributions. In terms of size measures of the confidence or credible sets, the Bayesian solutions are vastly superior, and between the two Bayesian credible sets, the HB-based sets are shorter or more “compact” than the UB-based sets.
Supplementary Materials
S.1 Gibbs sampling: The set of full conditional pdfs for the Gibbs sampling.
S.2 Size measures of elliptical sets: Discussion of two size measures of elliptical sets.
S.3 More results on commuting and income data: A double scatter plot, Figure S.1, from the HB posterior sample of commuting times discussed in Section 8, an extraction of the double stochastic matrix, Table S.1, from the same posterior sample, and Table S.2, a table of average or expected absolute deviations of ranking from the surrogate median incomes discussed in Section 9.
S.4 Ranking of baseball player: A full discussion of the application ranking of American major league professional baseball batters, discussed briefly in Section 10. This example was considered by Efron and Morris (1975) demonstrating the usefulness of EB inference.
Disclosure Statement: No competing interests to declare from the authors.
Funding
This work was partially supported by the National Science Foundation (NSF) grant DMS-2311186.
Supplementary Materials
S.1 The set of full conditional pdfs for the Gibbs sampling
The joint posterior pdf of $\theta,\beta,A$ is
$\pi(\theta,\beta,A|y)$ $\propto$ $\exp[-\frac{1}{2}(y-\theta)^{T}D^{-1}(y-\theta)]\times A^{-\frac{m}{2}}\exp[-\frac{1}{2A}(\theta-X\beta)^{T}(\theta-X\beta)]$
$\times$ $(\bar{D}+A)^{-\frac{1}{2}}.$
From the above the full conditionals are
  • (A)
    $\theta|\beta,A,y\sim N(\delta,\Phi),$
  • (B)
    $\beta|\theta,A,y\sim N((X^{T}X)^{-1}X^{T}\theta,A(X^{T}X)^{-1})$,
  • (C)
    $\pi(A|\theta,\beta,y)\propto(\bar{D}+A)^{-\frac{1}{2}}A^{-\frac{m}{2}}\exp[-\frac{1}{2A}(\theta-X\beta)^{T}(\theta-X\beta)]$,
where
$\Phi=(D^{-1}+A^{-1}I_{m})^{-1},~{}\delta=\Phi(D^{-1}y+A^{-1}X\beta),$
To sample from the conditional pdf of $A$ we use the rejection sampling. We use the inverse gamma distribution with shape parameter $(m-1)/2$ and rate parameter $(\theta-X\beta)^{T}(\theta-X\beta)/2$ as the proposal distribution. Alternatively, one can also use the Metropolis-Hastings algorithm within Gibbs with the same proposal as the transition density for $A$.
S.2 Size measures of elliptical sets
As an alternative to an orthotope (or hyperrectangular) credible set, where each component of $\theta$ varies freely, we have considered HPD sets which are approximately elliptical for our Bayesian models. Two size measures that appear to be appropriate for set estimators are their Lebesgue volumes and average lengths of the “sides” determining the set. These two measures are easily found for an orthotope. We compute representative values of these size measures of the elliptical sets.
To determine reasonable size measures for elliptical regions, we proceed with the set
$E=\{x:x^{T}\Phi x\leq c\},$ (S.1)
where $\Phi$ is an $m\times m$, p.d. matrix, and $c>0$, determine the region. Since the center of the ellipse has no role in the size measures, we take it zero. Let $V_{E}(\Phi,m,c)$ denote the Lebesgue volume of the set $E$. It can be easily calculated that
$V_{E}(\Phi,m,c)=\frac{\pi^{m/2}}{\Gamma((m+2)/2)}\times\frac{c^{m/2}}{\mbox{det}(\Phi)^{1/2}},$ (S.2)
where $\Gamma(\cdot)$ is the gamma function and $\mbox{det}(\Phi)$ is the determinant of $\Phi$.
A standard length measure of an elliptical region is not available. We propose below a reasonable expression for it. For an $m$-dimensional rectangular region, its volume is the product of the volume of the $(m-1)$-dimensional of the cross-section of the set defined by the intersecting hyperplane, say, $\theta_{1}=d$, and the length $(U_{1}-L_{1})$ of original rectangular set along the $\theta_{1}$ direction (where $L_{1}<d<U_{1}$). Thus, the length of the set along the $x_{1}$ direction is the ratio of the volume of the original orthotope to the volume of the lower-dimensional cross-section orthotope.
We will use the above discussion for the orthotope region as an analogy to come up with a suitable “length” for the ellipse along the $x_{1}$ direction. For the ellipse, we take for any $x_{1}$ and $x_{(2)}=(x_{2},\cdots,x_{m})^{T}$ with $(x_{1},x_{(2)}^{T})^{T}$ inside the set, we define a central cross-section by the hyperplane $x_{1}=-\Phi_{12}x_{(2)}/\Phi_{11}$, where we partition the matrix $\Phi$ as $((\Phi_{ij}))$, $i,j=1,2$, in conformity with the partitioning of $x$ above. We also define the p.d. matrix $\Phi_{22.1}=\Phi_{22}-\Phi_{21}\Phi_{11}^{-1}\Phi_{12}$. The corresponding cross-section is $E_{C}=\{x_{(2)}^{T}\Phi_{22.1}x_{(2)}\leq c\}.$ From equation (S.2), it similarly follows that
$V_{E_{C}}(\Phi_{22.1},m-1,c)=\frac{\pi^{(m-1)/2}}{\Gamma((m+1)/2)}\times\frac{c^{(m-1)/2}}{\mbox{det}(\Phi_{22.1})^{1/2}}.$ (S.3)
A reasonably representative length for the $x_{1}$ side of the ellipse is the ratio of the volume in (S.2) to the volume in (S.3). Simplification leads to an expression. From this, after simplification, in general, we have along the $x_{i}$ direction the representative length, $L_{R,i}$ (say), is
$L_{R,i}=B\left(\frac{1}{2},\frac{m+1}{2}\right)\times\left\{\frac{\sqrt{c}}{\sqrt{\Phi_{ii}}}\right\},~{}i=1,\cdots,m,$ (S.4)
where $B(a,b)$ is the regular beta function.
The above proposal of the length of a side of an elliptical region is only heuristic. Indeed it is a reasonable requirement that these lengths should be “calibrated” by a suitable scale so that for the modified orthotope based on the lengths of the “scaled sides” will match the volume of the original ellipse. Based on this consideration, the modified lengths $L_{M,i}$ are given by
$L_{M,i}=\{V_{E}(\Phi,m,c)/\prod_{i=1}^{m}L_{R,i}\}^{1/m}L_{R,i}.$ (S.5)
$L_{E}\stackrel{{\scriptstyle def}}{{=}}\frac{1}{m}\sum_{i=1}^{m}L_{M,i}$, a “linear” measure of the elliptical set, may be compared to $L_{C}$.
S.3 More Analysis for Commuting and Income Data
[…]
Table S.1: An extraction of a $9\times 9$ double stochastic matrix for the CD from the HB HPD (weighted) credible set for the commuting example
States
Rank ID KS IA AK MT NE WY ND SD
1 0.00 0.00 0.00 0.00 0.00 0.00 0.01 41.92 58.07
2 0.00 0.00 0.00 0.00 0.00 0.00 0.11 57.96 41.92
3 0.00 0.00 0.00 8.78 20.77 31.76 38.57 0.11 0.01
4 0.00 0.00 0.00 11.37 20.78 49.58 18.27 0.00 0.00
5 0.00 0.11 0.60 22.40 41.86 14.64 20.40 0.00 0.00
6 0.00 2.43 7.06 49.37 16.30 4.02 20.82 0.00 0.00
7 0.00 31.77 61.13 5.92 0.26 0.00 0.92 0.00 0.00
8 0.10 65.61 31.21 2.16 0.04 0.00 0.89 0.00 0.00
9 99.90 0.08 0.00 0.00 0.00 0.00 0.02 0.00 0.00
Table S.2: A table of average or expected absolute deviations of ranking from the surrogate median income data from 1990 census
KWW HB-Ellip UB-Ellip
MS 22.50 2.87 5.76
AR 22.54 2.63 8.45
NM 20.13 1.97 2.05
WV 19.76 1.63 3.25
LA 20.41 5.95 12.65
SD 18.64 2.78 3.67
KY 18.38 3.50 3.95
MT 15.77 3.30 4.92
OK 18.41 4.70 19.75
ID 15.91 2.91 3.44
ND 16.24 3.42 7.38
AL 15.25 2.66 5.96
TN 14.32 3.08 7.13
SC 14.29 2.21 6.16
TX 12.27 3.47 6.31
NE 14.71 5.52 13.00
UT 13.94 2.58 6.09
WY 14.00 6.13 6.79
IA 12.98 3.30 4.64
NC 12.42 2.68 2.23
AZ 13.24 4.03 9.30
OR 13.06 3.85 9.33
MO 12.92 4.05 7.57
KS 12.82 4.93 11.83
ME 12.50 4.97 5.82
FL 12.04 3.60 7.10
KWW HB-Ellip UB-Ellip
GA 12.76 4.98 15.00
VT 12.82 3.80 6.51
IN 12.92 3.43 5.33
OH 12.12 5.70 10.69
WI 13.24 3.27 5.05
PA 13.24 3.31 4.15
CO 13.55 5.08 14.80
NV 13.90 5.09 21.42
WA 14.33 2.91 5.48
MI 13.90 3.89 3.29
MN 15.12 3.44 5.53
VA 15.57 5.54 8.41
DC 16.06 5.99 17.91
IL 16.70 2.17 5.43
RI 17.16 3.09 4.06
NY 15.60 1.76 4.40
CA 18.86 4.96 12.82
NH 15.30 2.39 4.04
DE 19.82 9.29 18.55
HI 20.59 3.05 7.57
MD 21.39 0.99 3.48
AK 22.24 1.28 15.31
MA 10.73 0.45 1.48
NJ 13.57 0.47 1.61
CT 18.00 0.68 1.22
Average 15.67 3.52 7.61
S.4 Ranking of baseball players: An example by Efron and Morris (1975)
Efron and Morris (1975) considered batting averages of 18 major league baseball batters from their first 45 at bats in the 1970 baseball season to predict their performances in the remainder of that season. Let $\theta_{i}$ be the unknown true batting average of the $i$th player during that season. A reasonable representative value for $\theta_{i}$ was the player’s average ($\omega_{i}$) in the remainder of the season after his first 45 at bats. The sample proportion of hits $Y_{i}$ from first 45 at bats is approximately normal with mean $\theta_{i}$ and variance $D_{i}$. In our analysis, we use the estimated variance, given by $D_{i}=Y_{i}(1-Y_{i})/45.$ Treating the available $\omega_{i}$ as the “gold standard” value of $\theta_{i}$, the total empirical squared error (TESE), $\sum_{i=1}^{18}(y_{i}-\omega_{i})^{2}$, for the direct estimates $y_{i}$’s, is $0.07537$, and the same error for the HB and the empirical best linear unbiased predictor (EBLUP) estimates of $\theta_{i}$’s are $0.02368$ and $0.02356$, respectively. The direct estimates have $3.18$ times larger TESE than the HB estimates, and $3.20$ times larger TESE than the EBLUP estimates.
Efron and Morris (1975) used the arc-sin transformed parameters $\psi_{i}=arcsin(2\theta_{i}-1)$, for the transformed data $\hat{\psi}_{i}=arcsin(2Y_{i}-1)$ in their EB analysis. This transformation resulted in equal approximate sampling variances, $1/45$, for all 18 $\hat{\psi}_{i}$’s. They applied Stein’s shrinkage estimation method based on an exchangeable model to calculate ${\hat{\psi}}_{J-S,i}$, the James-Stein estimate of $\psi_{i}$. The TESE of prediction for the direct estimates $\hat{\psi}_{i}$ from $arcsin(2\omega_{i}-1)$ is 0.3902, which is about 3.5 times of 0.1113, the TESE for the James-Stein estimates ${\hat{\psi}}_{J-S,i}$. This application showed a substantial efficiency of Stein’s method. The efficiency for HB estimators of $\psi_{i}$ based on the same model is about $3.36$, which is comparable. Note that these numbers are comparable to $3.18$ and $3.20$, we obtained above for EBLUP and HB predictors of $\theta_{i}$’s without transformation.
Figure S.2: An alternative visualization of the Klein et al. (2020) 90% joint confidence region for overall ranking of baseball players with overlapping CDs for UB and HB models (weighted, elliptical method)
By equation (9) of Klein et al. (2020) a $90\%$ joint confidence region for the rank vector $\check{r}$ is given by
$\{(\check{r}_{1},\cdots,\check{r}_{18}):\check{r}_{k}\in\{1,2,\cdots,18\}\mbox{ for }k=1,\cdots,18\}.$
This confidence set for the rank vector is depicted in Figure S.2. Possible ranks for each player in this set are shown by a yellow line segment stretching from 1 to 18. According to this confidence set, any player can rank from 1 to 18, and any particular rank can be held by any of these 18 players; clearly, it is totally uninformative. Due to multiple testing, the critical value $2.76$ (cf. equation (13) of Klein et al. (2020)) is too large to declare any comparison between two direct estimates to be significant. A joint $90\%$ credible distribution for the rank vector is a double-stochastic matrix, each row and each column of the matrix sums to one. This probability matrix is pictorially presented by overlaying on the figure of the confidence set solid red circles with area of a circle is proportional to the probability it is representing. The top matrix is based on the elliptic set for $\theta$ under the UB model and the bottom matrix is a similar version for the intercept only HB model. From the figure, it is clear that many of the probabilities in this matrix are practically zeros implying that those combinations of players and ranks are improbable. Undoubtedly, the credible distributions are lot more informative than the confidence set.
The observed rank $R_{i}$ of the $i$th player based on $y$ is shown by a blue square on the $R_{i}$th row and the $i$th column of the figure. If several players tie, we follow the recommendation of Klein et al. (2020) to assign the highest possible rank to the tied players. We used $\omega_{i}$ values to find the surrogate rank vector $\xi$ for the true vector $\check{r}$. The $i$th component, $\xi_{i}$, for player $i$ is similarly depicted by a black circle on the $(\xi_{i},i)$ cell of the matrix.
The approach to the solution in the top panel of Figure S.2 is the one which is the most similar to the Klein et al. (2020) solution. It is clear from the first row of this panel is that Player 1 has the largest probability of being the best hitter, and only the first five or six players have realistic chances to be the best hitter. Similarly, some of the players 12 through 18 are the most likely ones who had the lowest performance. These lists of players are much narrower than the whole set of $18$ players suggested by Klein et al. (2020) method. From the first column of this panel, on the other hand, we find that a realistic list of ranks for Player 1 is $\{14,15,16,17,18\}$. This list is much narrower than the entire set $\{1,2,\cdots,18\}$ suggested by the frequentist method. This figure also indicates that most likely ranks for player 17 are 1 to 7. The rank $17$ is most likely to be held by players 1 to 7. Clearly, the Bayesian solution is substantially more informative.
In the bottom panel of Figure S.2 we depicted the double stochastic matrix from the HPD credible set for $\theta$ using the HB model. Due to shrinkage of the joint posterior distribution towards a vector of common means, probability distributions of the rankings of the few uppermost and the lowermost players are not as sharp as those provided by the corresponding credible distributions from the UB model. Even though there is a weakness of the HB method for its reliance on shrinkage to a common mean in the absence of a useful covariate, the method provides better solutions than the UB method for some players whose early batting averages are at odds with their latter scores. Here, five players (4, 8, 13, 16, 17) showed a large or moderately large discordance between their early and late season performances. Two most discordant players are 4 and 17. Player 4 did not perform well in the remainder of the season which earned him a very low surrogate rank $3$ compared to his high $y$ rank $15$. From the top panel of Figure S.2 we see that whereas the UB method assigns, perhaps erroneously, nearly 87% probability of Player 4 having a rank higher than 9, the bottom panel of Figure S.2 reveals that the HB method correctly assigns less than 63% probability to this player having such a high rank. At the other end, Player 17 performed very well in the remainder of the season which earned him a very high surrogate rank $17$ (vs. a very low $y$-rank 2). Here, again, Figure S.2 for the unweighted UB Cartesian method assigns, perhaps erroneously, nearly 93% probability of it having a rank lower than 10, the bottom panel of Figure S.2 for the HB method correctly assigns less than 67% probability to it having such a low rank. Revised rankings of the discordant players afforded by the HB method are noteworthy and more reliable. We have demonstrated through simulation in Subsection 10.1 and the median income example in Section 9 that the HB method in the presence of useful covariates is substantially more effective.
Also, for each player the variance of the possible ranks in the joint Klein et al. (2020) confidence region is $26.9$, the largest possible value, and the corresponding posterior variances from the credible distributions of ranks for these players vary between 5.36 and 16.36 for the UB model, and between 16.75 and 23.40 for the HB model. These numbers show a superiority of the credible distributions, created from the UB model or the HB model. We provide additional comparisons of the frequentist and the Bayesian solutions in Table  S.3.
Table S.3: Batting averages of 18 Major League players in 1970: $y_{i}$ is the average from first 45 at bats, $\omega_{i}$ the remainder of the season average, $R_{i}$ rank of $y_{i}$, $\xi_{i}$ rank of $\omega_{i}$, and deviation measures described in equation (14) for five overall ranking set estimators. For each of the two Bayesian models, we have two methods based on elliptical (subscript E) and Cartesian (subscript C) sets. We used the weighted method here.
Players $y_{i}$ $\omega_{i}$ $R_{i}$ $\xi_{i}$ $\epsilon_{i,\small{HB,C}}$ $\epsilon_{i,HB,E}$ $\epsilon_{i,UB,C}$ $\epsilon_{i,UB,E}$ $\epsilon_{i,KWW}$
1. Clemente .400 .346 18 18 5.90 5.90 1.18 1.18 8.50
2. F. Robinson .378 .298 17 15 4.39 4.39 1.87 1.87 6.17
3. F. Howard .356 .276 16 13 4.24 21.0 2.73 2.73 5.17
4. Johnstone .333 .222 15 3 8.11 8.11 11.40 11.39 6.83
5. Berry .311 .273 13.5 12 4.33 4.33 2.44 2.44 4.83
6. Spencer .311 .270 13.5 11 4.25 4.25 2.96 2.96 4.61
7. Kessinger .289 .263 12 7 4.96 4.96 4.92 4.92 4.83
8. L. Alvarado .267 .210 11 2 7.66 7.66 8.13 8.13 7.61
9. Santo .244 .269 9.5 10 4.36 4.34 3.09 3.09 4.50
10. Swoboda .244 .230 9.5 5 5.41 5.41 3.98 3.98 5.61
11. Unser .222 .264 6 8.5 4.26 4.26 3.09 3.08 4.56
12. Williams .222 .256 6 6 4.40 4.40 2.64 2.64 5.17
13. Scott .222 .303 6 16 7.64 7.64 9.36 9.36 6.83
14. Petrocelli .222 .264 6 8.5 4.25 4.25 3.33 3.33 4.56
15. E. Rodriguez .222 .226 6 4 5.38 5.38 3.46 3.46 6.17
16. Campenaris .200 .285 3 14 6.64 6.64 8.97 8.97 5.61
17. Munson .178 .316 2 17 9.54 9.54 13.57 13.56 7.61
18. Alvis .156 .200 1 1 5.46 5.46 1.43 1.43 8.50
Total abs. dev. ($\check{r}$) 101.16 101.16 88.56 88.56 107.64
$10^{13}\times$ Vol. ($\theta$) 51.38 1.32 93756 1757 85722
Average length ($\theta$) 0.236 0.193 0.359 0.288 0.357
From Table S.3 we find that the Klein et al. (2020) method has the highest sum of average absolute deviations (based on equation (14), about 6.4% larger than for the HB method, and about 21.5% larger than for the UB method. This is counter-intuitive as the shrinkage estimators of $\theta_{i}$’s from the exchangeable HB model have lower total squared error than the same for the direct estimators $Y_{i}$’s, which are Bayes estimates of $\theta_{i}$ under the UB model. We have noted above that the total squared error of UB estimate of $\theta$ from the “true” vector $\omega$ was $0.07537$, and for the HB estimate it was $0.02368$. We present below a possible explanation for the domination of the UB method over the HB method when it comes to inference for overall ranking based on their credible distributions.
In this example, estimating $\theta_{i}$ from the HB model shrinks the direct estimate; the shrinkage of $y_{i}$ towards a grand mean, measured roughly by the posterior mean of $E[D_{i}/(D_{i}+A)|y]$, is substantial (ranging between $0.75$ and $0.63$, and the shrinkage is more for the larger $D_{i}$’s that correspond to larger $y_{i}$’s. In our example, $D_{i}$’s decrease from $0.0053$ to $0.0029$, and the posterior mean of $A$ is $0.0024$ and the median of $A$ is $0.0017$.) These indicate that even in comparison with the small $D_{i}$’s, $A$ is small. In the absence of a covariate, a very small model error variance for an intercept only model essentially forces that $\theta_{i}$’s are all nearly equal. Accurate inference for the rank vector for such $\theta$ will naturally be challenging. Among the components of $\theta$, any differentiating information about $\theta_{i}$’s is given by its individual $y_{i}$’s only. Clearly, any shrinkage of an estimator of $\theta_{i}$ to a common value will reduce the estimator’s reliance on individual $y_{i}$’s, and thus the HB posterior distribution of $\theta_{i}$’s for ranking will be less informative than $y_{i}$, which is the center of the individual UB posterior distributions.
Due to lack of a covariate, corresponding to the set of larger observations (the first 4 or 5 observations which are in descending order), the HB posterior distribution of any of these $\theta_{i}$’s will shrink downward (to some grand mean). This downward shrinking of the posterior distribution of $\theta_{i}$’s will lower the probability of them holding the higher ranks (18 through 15 or 14). In fact, from Table S.4 the posterior expectation of $\check{r}_{1}$ in our example is $12.1$, which is quite a bit smaller than the largest rank $18$ (in this case) it is supposedly estimating. Similarly, for the next two smaller observations, the posterior expected ranks are again quite smaller than the target ranks 17 and 16. These discrepancies contribute to a bigger absolute deviations reported in Table S.3. On the other hand, the distributions of $\check{r}_{1},\check{r}_{2},\check{r}_{3}$ for the UB model do not shrink downward much since the posterior of $\theta_{i}$’s are independent and each is concentrated around its $y_{i}$. The corresponding posterior means of $\check{r}_{1},\check{r}_{2},\check{r}_{3}$ are $16.8,16.2$ and $15.2$ (cf. Table S.4), which are not too far from $18,17$ and $16$, respectively. These lesser discrepancies make smaller contributions to the expected absolute deviations reported Table S.3. A similar shrinkage upward will be noticed for the set of smaller observations (observations 14 through 18). Corresponding expected absolute deviations of the rank from the surrogate true rank $\xi_{i}$ for observations 14 through 18 under the HB model again tend to be bigger than their counterparts from the UB model. We see bigger discrepancies between the HB ranks and the UB ranks for the first 4 and the last 4 observations. All of these mostly explain the smaller total posterior expected absolute rank deviations for the UB model in comparison with their counterparts for the HB model.
Table S.4: Estimation of ranks of baseball players by different methods (weighted versions were used for Bayesian methods)
Players P01 P02 P03 P04 P05 P06 P07 P08 P09
$y$ ranks 18 17 16 15 14 13 12 11 10
$\omega$ ranks 18 15 13 3 12 11 7 2 10
KWW 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50
HB-Cart 12.10 11.86 11.60 10.94 10.41 10.94 10.27 9.57 9.35
HB-Ellip 12.10 11.86 11.60 10.94 10.41 10.94 10.27 9.57 9.35
UB-Cart 16.82 16.20 15.23 14.40 12.98 12.80 11.51 10.12 8.40
UB-Ellip 16.82 16.20 15.23 14.39 12.98 12.79 11.51 10.12 8.40
Players P10 P11 P12 P13 P14 P15 P16 P17 P18 Average
$y$ ranks 9 8 7 6 5 4 3 2 1 9.5
$\omega$ ranks 5 8.5 6 16 8.5 4 14 17 1 9.5
KWW 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.5
HB-Cart 9.22 8.61 8.54 8.60 8.56 8.56 7.91 7.50 6.46 9.5
HB-Ellip 9.22 8.61 8.54 8.60 8.56 8.56 7.92 7.50 6.46 9.5
UB-Cart 8.41 6.71 6.69 6.64 6.43 6.79 5.03 3.43 2.43 9.5
UB-Ellip 8.41 6.72 6.69 6.64 6.43 6.79 5.03 3.44 2.43 9.5
Even though the UB method is poised to have smaller total expected absolute deviations than the HB method without a covariate, for some individual players the HB method is better suited to estimate ranks than the UB method. Continuing the above discussion further, since the posterior distribution of $\theta_{i}$ under the UB model is determined only by $y_{i}$, in case of a big discrepancy between $y_{i}$ and the unknown $\theta_{i}$ it is estimating (which will be reflected through a bigger difference between $y_{i}$ and $\omega_{i}$, the surrogate value of $\theta_{i}$), the posterior expected absolute deviation of $\check{r}_{i}$ from $\xi_{i}$ will be large ($\check{r}_{i}$, the rank of $\theta_{i}$, and $\xi_{i}$, the rank of gold standard value $\omega_{i}$). Indeed this happens for five players 4, 8, 13, 16 and 17 with large discordance between a player’s early and late season performances. Among these five players, player 17 has the second smallest observation, who displayed large discrepancies between $y_{i}$ and $\omega_{i}$. Deviations $\epsilon_{i,UB}$ for these players are much larger than the average deviations of the other players, and these deviations also exceed the corresponding players’ deviations $\epsilon_{i,HB}$ based on the exchangeable HB model. If one or more player substantially over-perform or under-perform in the remainder of the reason, based on their later performances the true ranks of these players will change substantially from their ranks determined by $y_{i}$’s. The HB method by virtue of its borrowing from the other players is better suited to rectify the erroneous ranks of the players based on $y_{i}$’s. This is the likely reason that the HB ranks for the five players identified above perform better than the UB ranks that are based on individual $y_{i}$ alone; a downward shrinkage of the HB posterior distribution of the over-performing (in their first 45 at bats) players 4 and 8, and an upward shrinkage of the HB posterior distributions of the under-performing (in their first 45 at bats) players 13, 16 and 17. The HB posterior distributions of $\theta_{i}$ of these players, by virtue of their borrowing information from the other players, are substantially pulled toward the “true value”, measured by $\omega_{i}$. This is a useful property of the HB model. No such borrowing of information is possible through the UB model as the posterior distributions of the components $\theta_{i}$’s are independent and centered around $y_{i}$.
Lastly, from this example where the total posterior expectation of absolute deviations of the true rank $\check{r}_{i}$ from the surrogate rank $\xi_{i}$ for the UB model is lower than the same for the HB model, we may not conclude that the UB model will always provide a more accurate solution. In fact, it is shown in simulations in the paper that if a covariate effectively explains the $\theta_{i}$’s, the credible distributions of overall ranking based on the HB model are superior to the credible distributions derived from the UB model. Intuitively, this occurs since the posterior distribution of $\check{r}_{i}$ under the HB model will better concentrate near the true rank than the same posterior distribution under the UB model. We should note that even though in terms of total of expected absolute deviations, the frequentist Klein et al. (2020) method fares the worst, it is surprisingly robust for the five under-/over-performing players 4, 8, 13, 16 and 17. Finally, we included some size measures of the confidence set (for the KWW method) and credible sets for Bayesian methods, the sets are critical for the ranking problem. The last two rows of Table S.3 provide volumes and approximate average lengths of the sets, respectively. In terms of volume, the HB sets are at least a thousand times better than the other sets. In terms of length, the HB sets are at least 50% better than the other sets. The best among all five sets is the HB elliptical set.